### v1

Abstract
Reinforcement learning agents can achieve superhuman performance, but their decisions are often difficult to interpret. This lack of transparency limits deployment, especially in safety-critical settings where human trust and accountability are essential. In this work, we develop a theoretical framework for explaining reinforcement learning through the influence of state features—what the agent observes in its environment. We identify three core elements of the agent-environment interaction that benefit from explanation: behaviour (what the agent does), performance (what the agent achieves), and value estimation (what the agent expects to achieve). We treat state features as players cooperating to produce each element and apply Shapley values—a principled method from cooperative game theory—to identify the influence of each feature. This approach yields a family of mathematically grounded explanations with clear semantics and theoretical guarantees. We use illustrative examples to show how these explanations align with human intuition and reveal novel insights. Our framework unifies and extends prior work, making explicit the assumptions behind existing approaches, and offers a principled foundation for more interpretable and trustworthy reinforcement learning.
Keywords: Reinforcement learning, Shapley values, explainable artificial intelligence, explainable reinforcement learning, Markov decision process, feature-based explanations
1 Introduction
Reinforcement learning is an approach to artificial intelligence that enables agents to learn through trial-and-error interaction with their environment. By maximising a scalar reward signal that provides feedback on their actions, agents learn to achieve desired outcomes. Reinforcement learning has achieved remarkable success across diverse domains, from mastering complex games (Mnih et al., 2015; Silver et al., 2016, 2017; Vinyals et al., 2019; Schrittwieser et al., 2020; Wurman et al., 2022) to controlling real-world systems (Bellemare et al., 2020; Degrave et al., 2022; Seo et al., 2024). Unlike humans, reinforcement learning agents typically cannot explain their behaviour, making it difficult for people to understand or trust their decisions. This hinders deployment, especially in safety-critical domains where human accountability is essential. For example, if a human operator could not understand a reinforcement learning agent’s decisions, would they deploy it to regulate a nuclear fusion reactor (Degrave et al., 2022; Seo et al., 2024)?
In this work, we address this limitation by introducing a theoretical framework for explaining reinforcement learning. We identify three core elements of reinforcement learning that benefit from explanation: behaviour (what the agent does), performance (what reward the agent achieves), and value estimation (what reward the agent expects to achieve). Together, these elements offer a comprehensive understanding of how agents interact with their environments.
We propose to explain these three elements through the influence of state features—numerical values that describe what the agent observes in its environment. By treating features as players cooperating to produce behaviour, performance, or value estimates, we cast the problem in terms of cooperative game theory. This framing allows us to attribute each feature’s contribution using Shapley values (Shapley, 1953)—a principled method for fairly distributing credit among contributors—with the resulting explanations uniquely characterised by a set of intuitive mathematical axioms. The outcome is a theoretically principled framework: Shapley Values for Explaining Reinforcement Learning (SVERL). Using illustrative examples, we show that SVERL produces explanations that align with human intuition while revealing novel insights into an agent’s behaviour, outcomes, and expectations.
Although Shapley values have been used in reinforcement learning (Rizzo et al., 2019; Carbone, 2020; He et al., 2021; Wang et al., 2020; Zhang et al., 2020; Liessner et al., 2021; Løver et al., 2021; Remman et al., 2022; Schreiber et al., 2021; Theumer et al., 2022; Zhang et al., 2021), prior work has largely focused on specific applications, offering no formal analysis of what the resulting explanations actually mean. Without a clear theoretical foundation, these methods leave key questions unanswered: What does a Shapley value explain in reinforcement learning? How should it be interpreted? And how do previous explanations relate to one another? In contrast, SVERL provides a principled mathematical foundation for addressing these questions. It situates existing approaches within a unified perspective, revealing their underlying assumptions and limitations, and showing how they can yield incomplete, misleading, or easily misinterpreted explanations.
The remainder of this paper is structured as follows. In Section 2, we provide useful background information on reinforcement learning and Shapley values. In Section 3, we introduce the motivation and foundational questions behind explaining reinforcement learning. In Sections 4, 5 and 6, we develop a unified framework for explaining agent behaviour, performance, and value estimation, respectively. In Section 7, we demonstrate the proposed framework in various illustrative domains. In Section 8, we discuss how the framework can be used, interpreted, and extended. We review related work in Section 9, propose directions for future research in Section 10, and conclude in Section 11.
2 Background
In this section, we introduce our notation and briefly describe reinforcement learning (Sutton and Barto, 2018), Shapley values (Shapley, 1953) and how Shapley values can be used to explain supervised learning models (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017; Frye et al., 2020a; Covert et al., 2021).
We indicate random variables by capital letters (e.g. $X$, $S$), their values by lowercase letters (e.g. $x$, $s$) and their domains with calligraphic fonts (e.g. ${\cal X}$, ${\cal S}$). Functions are also denoted by lowercase letters (e.g. $g$); it will be clear from context whether a lowercase letter refers to a function or a value. Random variables are assumed to be vectors, with ${\cal X}_{i}$ indicating the domain for value $x_{i}$ at index $i$. Given a base set =${\cal F}=\{1,2,\dots,n\}$, where $n$ is the size of the domain ${\cal X}$, we use $\bar{\mathcal{C}}\equiv{\cal F}\setminus\mathcal{C}$ to represent a set’s complement, and $x_{\mathcal{C}}\equiv\{x_{i}:i\in\mathcal{C}\}$ to denote a subset of values for $\mathcal{C}\subseteq{\cal F}$. Additionally, we use $\Delta({\cal A})$ to denote the probability simplex over the domain ${\cal A}$. That is, the function $\rho:S\rightarrow\Delta({\cal A})$ expresses a probability mass function $\rho(\cdot|s)$ over ${\cal A}$, for each $s\in{\cal S}$.
2.1 Reinforcement Learning
Reinforcement learning is an approach to artificial intelligence that models the interaction of an agent with its environment as the agent seeks to achieve desired outcomes. We follow the common approach of modelling this interaction as a Markov Decision Process (MDP), defined as a 5-tuple $({\cal S},{\cal A},p,r,\gamma)$. The environment is initiated in state $S_{0}\in{\cal S}$, according to some initial state distribution $d(s)\coloneqq\text{Pr}(S_{0}=s)$. At decision step $t$, $t\geq 0$, the agent observes the current state of the environment, $S_{t}\in{\cal S}$, and executes action $A_{t}\in{\cal A}(s_{t})$, where ${\cal A}(s_{t})$ is the set of executable actions in $s_{t}$. Consequently, the environment transitions to a new state, $S_{t+1}\in{\cal S}$, according to the transition kernel $p(s^{\prime}\,|\,s,a)\coloneqq\text{Pr}(S_{t+1}=s^{\prime}\,|\,S_{t}=s,A_{t}=a)$, and returns a reward $R_{t+1}\in\mathbb{R}$, whose expectation for a transition from state $s$ to state $s^{\prime}$ under action $a$ is denoted $r(s,a,s^{\prime})=\mathbb{E}[R_{t+1}\,|\,S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime}]$. The return $G_{t}$ from decision step $t$ is defined as the discounted sum of all future rewards:
$G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1},$
where $\gamma\in[0,1]$ is the discount factor.
We assume that states can be decomposed into features, $S=(S_{1},S_{2},\dots,S_{n})$, where each $S_{i}\in{\cal S}_{i}$ represents the value of an individual feature, such as the velocity of an autonomous vehicle. We will use these state features as the basic elements of our explanations of reinforcement learning.
A policy $\pi:{\cal S}\rightarrow\Delta({\cal A})$ maps each state to a probability distribution over actions. The agent’s objective is to learn an optimal policy that maximises the expected return from every state. Reinforcement learning methods commonly compute the optimal policy by alternating between policy evaluation: estimating the expected return from state $s$ when following policy $\pi$, $v^{\pi}(s)\coloneqq\mathbb{E}_{\pi}{\left[G_{t}\,|\,S_{t}=s\right]}$, or the expected return of executing action $a$, $q^{\pi}(s,a)\coloneqq\mathbb{E}_{\pi}\left[G_{t}\,|\,S_{t}=s,A_{t}=a\right]$, and policy improvement: deriving a new policy $\pi^{\prime}$ by acting greedily with respect to $q^{\pi}(s,a)$:
$\pi^{\prime}(s)\coloneqq\underset{a\in{\cal A}}{\operatorname*{arg\,max}}\,q^{\pi}(s,a).$
An elegant and widely-used algorithm for learning the value of the optimal policy is Q-Learning (Watkins and Dayan, 1992). With this algorithm, $q^{\pi}(s,a)$ is estimated by some representation $Q(s,a)$, which is updated using the temporal-difference update rule below, where $\alpha$ is the step-size parameter:
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha\left(R_{t+1}+\gamma\max_{a\in{\cal A}}Q(S_{t+1},a)-Q(S_{t},A_{t})\right).$
Reinforcement learning agents can use algorithms like Q-learning to learn an optimal policy. However, they do not explain the reasoning behind their behaviour while learning or executing the policy.
2.2 Shapley Values in Game Theory
When individuals collaborate to achieve a shared outcome, a fundamental question arises: How should the outcome be fairly divided among them? Cooperative game theory provides a framework for addressing this question by modelling player interactions. This framework produced the Shapley value, introduced by Shapley (1953) to fairly assign credit to each participant based on their contribution in different possible coalitions.
Formally, a coalitional game consists of a set of players, ${\cal F}$, and a characteristic function $v:2^{|{\cal F}|}\to\mathbb{R}$, which assigns a value to each coalition $\mathcal{C}\subseteq{\cal F}$, representing the outcome they achieve together. That is, $v(\mathcal{C})$ denotes the payoff obtained by the subset of players $\mathcal{C}$. By convention, the value assigned to an empty coalition, $v(\emptyset)$, is often set to zero: $v(\emptyset)=0$ (von Neumann and Morgenstern, 1947), though in some cases, a fixed baseline value is used instead (Aas et al., 2021). Coalitional games are sometimes also assumed to satisfy superadditivity, meaning that merging coalitions should not decrease their total payoff:
$v(\mathcal{C}\cup\mathcal{D})\geq v(\mathcal{C})+v(\mathcal{D}),\quad\forall\mathcal{C}\subseteq{\cal F},\forall\mathcal{D}\subseteq{\cal F}\text{ such that }\mathcal{C}\cap\mathcal{D}=\emptyset.$
We do not assume superadditivity because it is not required for deriving Shapley values and is incompatible with how they are applied in this work.
As an example of a coalitional game, consider a parliament with 100 members, where a vote requires a simple majority to pass; parties A and B both have 49 representatives, and party C has 2 representatives. With each representative given one vote, assumed to agree with their party, we can formulate a vote as a coalitional game played by the three political parties. The outcome of the game is measured by a characteristic function that returns 0 if the vote fails and 1 if the vote passes: $v(\emptyset)=0$, $v(\{\text{A}\})=0$, $v(\{\text{B}\})=0$, $v(\{\text{C}\})=0$, $v(\{\text{A},\text{B}\})=1$, $v(\{\text{A},\text{C}\})=1$, $v(\{\text{B},\text{C}\})=1$ and $v(\{\text{A},\text{B},\text{C}\})=1$.
A coalitional game allows us to investigate contribution schemes. For example, we might ask, how much does party A contribute to the outcome of a vote relative to parties B and C? Shapley values (Shapley, 1953) provide a principled approach to answering such questions by evaluating how the absence of each player affects a game’s outcome. The Shapley value of player $i\in{\cal F}$ in coalitional game $({\cal F},v)$, denoted by $\phi_{i}\left(v\right)$, is given by:
$\phi_{i}\left(v\right)=\sum_{\mathcal{C}\subseteq{\cal F}\setminus\left\{i\right\}}{\frac{\left|\mathcal{C}\right|!\leavevmode\nobreak\ \left(\left|{\cal F}\right|-\left|\mathcal{C}\right|-1\right)!}{\left|{\cal F}\right|!}\leavevmode\nobreak\ \big{(}v\left(\mathcal{C}\cup\left\{i\right\}\right)-v\left(\mathcal{C}\right)\big{)}}.$ (1)
Whilst the definition of a coalitional game $({\cal F},v)$ does not impose an ordering over players, introducing one provides an intuitive interpretation of the Shapley value. Imagine the players in ${\cal F}$ joining the game sequentially. When player $i$ joins an existing coalition $\mathcal{C}$, their contribution is the marginal gain in characteristic value: $v\left(\mathcal{C}\cup\left\{i\right\}\right)-v\left(\mathcal{C}\right)$. The Shapley value evaluates this marginal gain for every possible coalition $\mathcal{C}$ that could form without player $i$, weighting each marginal gain by a multinomial coefficient to account for every possible order players in $\mathcal{C}$ may join before player $i$ and players in ${\cal F}\setminus\mathcal{C}$ may join after player $i$. Therefore, an equivalent formulation expresses the Shapley value as the expected marginal contribution of $i$ over all possible orderings of players in ${\cal F}$:
$\phi_{i}\left(v\right)=\underset{O\sim\pi({\cal F})}{\mathbb{E}}\left[v\left(b^{i}(O)\cup\{i\}\right)-v\left(b^{i}(O)\right)\right],$ (2)
where $\pi({\cal F})$ is the set of all permutations of ${\cal F}$, and $b^{i}(O)$ represents the set of players that precede $i$ in ordering $O$.
In the parliamentary voting example, the Shapley values for the three parties are: $\phi_{\text{A}}=\frac{1}{3}$, $\phi_{\text{B}}=\frac{1}{3}$, $\phi_{\text{C}}=\frac{1}{3}$. Using the Shapley values, we can conclude that all three parties make the same contribution towards each vote and hence have the same voting power.
Shapley values are the unique solution to four axioms, described below, that specify the conditions that must be met for a fair allocation of contributions among the players of a coalitional game. While there are other contribution schemes in the literature, Shapley values offer the only contribution scheme derived from these axioms of fairness, accounting for interdependencies between players (often observed in practice).
Let $\phi_{i}\left(v\right)$ denote how much player $i$ contributes to the outcome of a coalitional game $({\cal F},v)$. For fair allocation of credit, $\phi_{i}\left(v\right)$, $\forall i$, must satisfy:
  • Axiom 1 (Efficiency). The difference in value between the outcome of the game, $v({\cal F})$, and the fixed payoff for the game with no players, $v(\emptyset)$, is fully distributed over the set of players:
    $v({\cal F})-v(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(v)$
  • Axiom 2 (Linearity). If two independent coalitional games, $u$ and $v$, are combined into a new game with the characteristic function $\alpha u+\beta v$ for some scalars $\alpha,\beta\in\mathbb{R}$, then each player’s contribution in the combined game must be the weighted sum of their contributions in the individual games:
    $\phi_{i}(\alpha u+\beta v)=\alpha\phi_{i}(u)+\beta\phi_{i}(v)$
  • Axiom 3 (Symmetry). Two players are assigned the same contribution if they make equal marginal contributions to the outcome of the game for all possible coalitions of players:
    $\phi_{i}(v)=\phi_{j}(v)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ v(\mathcal{C}\cup\{i\})=v(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\,\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A player is assigned zero contribution if they make zero marginal contribution to the outcome of the game for all possible coalitions of players:
    $\phi_{i}(v)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ v(\mathcal{C}\cup\{i\})=v(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\,\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
The seminal paper by Shapley (1953) proved that Shapley values provide the unique solution that satisfies these four axioms. [^1] The efficiency, linearity, and symmetry axioms are sufficient to derive the Shapley value, while the nullity axiom is a direct consequence of them.
2.3 Explaining Supervised Learning with Shapley Values
Shapley values have been adopted in supervised machine learning to explain model loss (Lipovetsky and Conklin, 2001; Owen, 2014; Lundberg et al., 2020; Covert et al., 2020; Williamson and Feng, 2020) and model predictions (Štrumbelj et al., 2009; Štrumbelj and Kononenko, 2010, 2011, 2014; Datta et al., 2016; Lundberg and Lee, 2017). In this section, we present an overview of the use of Shapley values to explain model predictions. For further details on using Shapley values to explain supervised machine learning, we refer the reader to the unified framework presented by Covert et al. (2021).
Consider a supervised learning model $f$ that predicts a target variable $Y\in{\cal Y}$ for an input $X=(X_{1},X_{2},\dots,X_{n})\in{\cal X}$, where $X_{i}\in{\cal X}_{i}$ represents the value of feature $i$. For example, $f$ could be a model that predicts the rating of a movie using the features genre, duration, and cast.
At the specific input point $x$, the prediction $f(x)$ is made collectively by the feature values $x_{1},x_{2},\dots,x_{n}$. This prediction can be modelled as a coalitional game $\text{G}=({\cal F},f_{x})$, where the players ${\cal F}$ are the feature values $x$. The outcome of the game is the prediction $f(x)$. The characteristic function is $f_{x}:2^{|{\cal F}|}\rightarrow{\cal Y}$, where $f_{x}(\mathcal{C})$ returns the prediction at $x$ when the values of features not in $\mathcal{C}\subseteq{\cal F}=\{1,2,\dots,n\}$ are unknown. The game quantifies how different subsets of feature values $(x_{1},x_{2},\dots,x_{n})$ contribute to the prediction $f(x)$; its Shapley values $\phi_{1}(f_{x}),\dots,\phi_{n}(f_{x})$ uniquely satisfy four axioms defining a fair credit allocation:
  • Axiom 1 (Efficiency). The difference between the prediction when the values of all features are known, $f_{x}({\cal F})$, and the fixed prediction when the values of no features are known, $f_{x}(\emptyset)$, is fully distributed over the feature values:
    $f_{x}({\cal F})-f_{x}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(f_{x})$
  • Axiom 2 (Linearity). A feature value’s contribution to the prediction of a linear sum of independent models, $f$ and $g$, is equal to the sum of the feature value’s contributions towards the predictions of each model:
    $\phi_{i}(\alpha f_{x}+\beta g_{x})=\alpha\phi_{i}(f_{x})+\beta\phi_{i}(g_{x})$
  • Axiom 3 (Symmetry). Two feature values that make equal marginal contributions to the prediction of a model, when used (for prediction) with all possible subsets of feature values, are assigned the same contribution:
    $\phi_{i}(f_{x})=\phi_{j}(f_{x})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{x}(\mathcal{C}\cup\{i\})=f_{x}(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature value that makes zero marginal contribution to the prediction of a model, when used with all possible subsets of feature values, is assigned zero contribution:
    $\phi_{i}(f_{x})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{x}(\mathcal{C}\cup\{i\})=f_{x}(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
The Shapley values attribute the change in prediction when all feature values are known versus when none are known, $f_{x}({\cal F})-f_{x}(\emptyset)$, to individual feature values. When a prediction $f(x)$ is multidimensional, such as the probability of belonging to different classes in a classification problem, independent Shapley values can be calculated for each element of $f(x)$.
Computing Shapley values requires making predictions $f_{x}(\mathcal{C})$ for all subsets of features $\mathcal{C}\subseteq{\cal F}$. However, the function $f(x):{\cal X}\rightarrow{\cal Y}$ specifies the prediction only when $\mathcal{C}={\cal F}$. It is therefore necessary to make some assumptions on how the model will predict when some features are absent. There are multiple approaches one can take. One common approach is to represent the unknown feature values by marginalising over their conditional distribution $p(X_{\bar{\mathcal{C}}}\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})$ (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017). Another common approach is to marginalise over their joint marginal distribution $p(X_{\bar{\mathcal{C}}})$ (Štrumbelj and Kononenko, 2010). We briefly explain these two approaches below.
The conditional characteristic function represents a model’s prediction when feature values $x_{\bar{\mathcal{C}}}$ are unknown by marginalising out the unknown feature values using their conditional distribution $p(X_{\bar{\mathcal{C}}}\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})$:
$f_{x}(\mathcal{C})=\mathbb{E}\left[f(X)\,|\,X_{\mathcal{C}}=x_{\mathcal{C}}\right].$
An in-depth discussion of the conditional characteristic function is given by Covert et al. (2020) and Covert et al. (2021), showing that conditional characteristic functions give rise to desirable information-theoretic properties. The marginal characteristic function, on the other hand, represents a model’s prediction when feature values $x_{\bar{\mathcal{C}}}$ are unknown by marginalising out the unknown feature values using their joint-marginal distribution $p(X_{\bar{\mathcal{C}}})$:
$f_{x}(\mathcal{C})=\mathbb{E}\left[f(x_{\mathcal{C}},X_{\bar{\mathcal{C}}})\right].$
As a consequence of assuming features are independent, marginal characteristic functions consider impossible feature value combinations when features are dependent (Frye et al., 2020a). They were once treated as a relatively easy approximation of conditional characteristic functions. However, it has been argued that marginal characteristic functions may be a better way to represent unknown feature values because the resulting Shapley value for a feature not explicitly used by a model is always zero; if using the conditional characteristic function, the Shapley value for such a feature may be non-zero (Sundararajan and Najmi, 2020; Janzing et al., 2020).
For example, the function $f(X_{1},X_{2})=X_{1}$ never uses the feature $X_{2}$. However, if $X_{2}$ depends on $X_{1}$ (or is correlated with $X_{1}$), using the conditional characteristic function results in non-zero Shapley values for $x_{2}$ at any point $x$. These Shapley values “reflect the model’s dependence on the information communicated by each feature [value], rather than its algebraic dependence” (Jethani et al., 2021). When using the marginal characteristic function, on the other hand, the Shapley values do not account for any relationship between $x_{1}$ and $x_{2}$ because they assume that the features are independent. In other words, with the marginal approach, the Shapley value for $x_{2}$ is zero: $\phi(x_{2})=0$.
In this paper, we mainly use conditional characteristic functions. However, the approach we propose is general and can be used with other ways of defining characteristic functions, including the marginal approach.
2.4 Global Explanations of Supervised Learning with Shapley Values
So far, we have used Shapley values to provide local explanations, which describe how individual feature values contribute to a single prediction $f(x)$. However, one may also want to understand how feature values influence predictions in general, leading to what are known as global explanations. While global Shapley-based explanations of model predictions have not been explored in the literature, similar methods have been used to explain a model’s loss.
Let $\ell(f(x),y)$ denote a loss function that quantifies the discrepancy between the model’s prediction $f(x)$ and the target variable $y$. Just as a single prediction can be framed as a coalitional game among feature values, Covert et al. (2020) propose framing a model’s expected loss as a coalitional game $\text{G}=({\cal F},v^{\ell})$, where the feature set ${\cal F}$ plays the game, and the characteristic function $v^{\ell}:2^{|{\cal F}|}\rightarrow\mathbb{R}$ assigns each subset $\mathcal{C}\subseteq{\cal F}$ the expected loss when the features outside $\mathcal{C}$ are unknown:
$v^{\ell}(\mathcal{C})=\mathbb{E}_{p(x)}\left[v^{\ell}_{x}(\mathcal{C})\right],$
which is an expectation of a per-instance characteristic function:
$v^{\ell}_{x}(\mathcal{C})\coloneqq\ell\left(f_{x}(\mathcal{C}),y\right).$
In this formulation, the global Shapley value of feature $i$, denoted $\phi_{i}(v^{\ell})$, represents its contribution to the overall prediction accuracy of the model. The characteristic function is a linear sum of per-instance characteristic functions $v^{\ell}_{x}(\mathcal{C})$ corresponding to separate coalitional games at each input $x$. Therefore, the linearity axiom ensures that the global Shapley values are also an expectation over the Shapley values for those coalitional games:
$\phi_{i}(v^{\ell})=\mathbb{E}_{p(x)}\left[\phi_{i}(v^{\ell}_{x})\right]$
This interpretation makes clear that a global Shapley value reflects the mean contribution of a feature to the model’s loss over the data distribution $p(x)$, providing insight into its overall importance across the dataset.
2.5 Computing Shapley Values in Practice
When the number of features or the domain ${\cal X}$ is large, the exact computation of Shapley values becomes intractable. In such cases, Shapley values and characteristic functions can be approximated. The conditional characteristic function can be approximated through Monte Carlo sampling (Štrumbelj and Kononenko, 2010, 2014):
$f_{\mathcal{C}}(x)=\mathbb{E}\left[f(X)\,|\,X_{\mathcal{C}}=x_{\mathcal{C}}\right]=\lim_{n\rightarrow\infty}{\frac{1}{n}\sum_{x^{\prime}\,\sim\,p(\cdot\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})}{f(x^{\prime})}},$ (3)
where the conditional distribution $p(X\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})$ can be approximated by a method such as variational inference (Frye et al., 2020a).
Two common methods use Equation 3 to approximate Shapley values in supervised learning. The first method combines Equation 3 and the Shapley values of Equation 2 into a single Monte Carlo sampling method (Štrumbelj and Kononenko, 2010, 2014):
$\phi_{i}(f_{x})=\lim_{n\rightarrow\infty}{\frac{1}{n}\sum_{\begin{subarray}{c}x^{\prime}\,\sim\,p(\cdot\,|\,X_{\mathcal{C}\cup\{i\}}=x_{\mathcal{C}\cup\{i\}})\\ \\ x^{\prime\prime}\,\sim\,p(\cdot\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})\end{subarray}}{f(x^{\prime})-f(x^{\prime\prime})}},$ (4)
where $\mathcal{C}$ is the set of features that precede $i$ in ordering $O$ and $O$ is sampled uniformly from $\pi({\cal F})$, the set of all permutations of ${\cal F}$. The second method is the kernel method, which approximates characteristic values using Equation 3, rewrites the Shapley values of Equation 1 as the solution to a particular weighted least squares problem, and then solves for the Shapley values for all features at a point $x$ simultaneously. The kernel method was introduced by Lundberg and Lee (2017); we direct the reader to Aas et al. (2021) for a detailed description of the approach.
This concludes our introduction to the relevant background. In the following section, we consider what aspects of agent-environment interaction can and should be explained (Section 3). We then formalise these explanations using Shapley values (Sections 4, 5 and 6).
3 Explaining Reinforcement Learning
A fundamental question in explaining reinforcement learning is: What exactly should be explained? Here, we propose one possible answer—explaining an agent’s interaction with its environment requires explaining three key elements:
  1. 1.
    Behaviour: How does the agent choose actions?
  2. 2.
    Performance: What reward do these actions earn?
  3. 3.
    Value estimation: How can we estimate these rewards?
Each of these elements provides a different perspective on the reinforcement learning problem. Together, they help us build a comprehensive understanding of an agent’s behaviour, the consequences of its decisions, and our ability to anticipate those consequences.
Consider an autonomous vehicle navigating a city using road signs (see Figure 1). The signs provide directional guidance and distance information, influencing the vehicle’s decision-making process. By analysing how the agent interprets these cues, selects actions, and achieves its goal, we can examine reinforcement learning through the following three perspectives.
Figure 1: An autonomous vehicle navigates toward a destination marked by flags, following road signs that provide directions and distances. The vehicle’s policy follows the black path, turning right at the first junction, while the red path represents an alternative shortest route.
  • Explaining Behaviour. An agent’s behaviour describes how it selects actions given its current knowledge. In our example (see Figure 1), the autonomous vehicle follows a policy that dictates its route based on road signs—such as turning right at the first junction because a sign indicates that direction. The vehicle might take a different route if a sign is missing or incorrect. Explaining behaviour means understanding why an agent makes certain decisions based on the available information.
  • Explaining Performance. Performance measures how well the agent achieves its objective, typically evaluated through the expected return of its policy. For instance, if the autonomous vehicle reaches its destination in minimal time, we consider it to have high performance. Performance is distinct from behaviour: a change in behaviour does not necessarily result in a change in performance. In Figure 1, the vehicle can turn either right or left at the first junction, but both paths lead to the destination in the same distance. While behaviour explanation concerns how an agent selects actions, performance explanation concerns how well those actions achieve the goal.
  • Explaining Value Prediction. Value estimation involves estimating how successful a behaviour will be. Consider the road sign distances in Figure 1. The first sign indicates that the destination is 10 miles away, while the second sign shows it is only 2 miles away. We expect a higher expected return at 2 miles than at 10 miles because the agent is closer to its goal. Importantly, estimating value considers how different inputs affect our estimate of expected return without necessarily changing the agent’s actual behaviour—the policy itself does not rely on the road sign distances. This distinguishes value estimation from behaviour and performance: the latter describes what actually happens, while the former describes what we expect to happen.
By explaining behaviour, performance, and value estimates, we gain a holistic understanding of reinforcement learning—covering an agent’s behaviour, the returns of that behaviour, and how we estimate those returns. In the following sections, we formalise these explanations using Shapley values. Specifically, we analyse how to explain behaviour (Section 4), performance (Section 5), and value estimation (Section 6), introducing our framework: Shapley Values for Explaining Reinforcement Learning (SVERL).
4 Explaining Agent Behaviour
To explain agent behaviour, we start by recognising that behaviour is determined by a policy: a function $\pi:{\cal S}\rightarrow\Delta({\cal A})$ that maps each state to a probability distribution over actions. We consider first discrete action spaces, before extending to continuous actions in Section 4.5. We assume that each state $s\in{\cal S}$ can be decomposed into the values of its features, $s=(s_{1},s_{2},\dots,s_{n})$. This common assumption is generally satisfied in practice. For example, an agent’s location can be represented by its $x$ and $y$ coordinates. Consequently, a sound approach to explaining agent behaviour is to analyse how a policy uses state features to determine $\pi(s,a)\in[0,1]$, the probability of selecting each action $a$ in state $s$, for all possible values of $s$ and $a$. For example, how does an agent’s $x-$coordinate influence the probability that it moves in a certain direction?
In the following sections, we present a theoretical analysis of explaining the behaviour of reinforcement learning agents using the influence of feature values on the probabilities of selecting actions.
4.1 State Features Collaborate to Determine Action Probabilities
Having established that agent behaviour can be understood through how feature values influence the probability of selecting actions, we now take a game-theoretic perspective on this relationship. Specifically, we observe that $\pi(s,a)$ can be viewed as the outcome of a coalitional game played by the features at state $s$, which collaborate to determine the probability of selecting action $a$. This naturally leads to the question: how much does each individual feature value contribute to the precise value of $\pi(s,a)$? To answer this, we define fairness axioms that any attribution method should satisfy and identify Shapley values as the unique solution. Below, we formally define the coalitional game, introduce the fairness axioms, and derive the unique solution from first principles.
Definition 1 (Discrete Policy Game)
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\tilde{\pi}_{s}^{a}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\tilde{\pi}_{s}^{a}(\mathcal{C})$ returns the probability of selecting action $a$ in state $s$ when the features $\mathcal{C}\subseteq{\cal F}$ are known by the agent and the features $\bar{\mathcal{C}}={\cal F}\setminus\mathcal{C}$ are unknown.
Central to the definition of the discrete policy game $({\cal F},\tilde{\pi}_{s}^{a})$ is the characteristic function $\tilde{\pi}_{s}^{a}:2^{|{\cal F}|}\rightarrow[0,1]$, which specifies how the agent would act—that is, the probability of selecting action $a$ when in state $s$—when only a subset of the feature values is known. This function depends on the underlying policy $\pi$, which governs behaviour when all state features are fully observed. However, the policy $\pi$ does not define the agent’s behaviour when some feature values are unknown. To account for this, we introduce the characteristic function $\tilde{\pi}_{s}^{a}$, which extends $\pi$ to settings where the agent has partial knowledge of the state. We use the same symbol $\pi$ to highlight this relationship: in the case where all features are known, that is, $\mathcal{C}={\cal F}$, we require that $\tilde{\pi}_{s}^{a}(\mathcal{C})=\pi(s,a)$. The precise construction of the characteristic function $\tilde{\pi}_{s}^{a}$ will be addressed later in Section 4.2. For now, we assume it is given and turn to the central question of allocating credit among the features.
The question we now face is: What is the precise contribution of each state feature value $s_{i}\in\mathcal{S}_{i}$ to the probability of selecting action $a$ in state $s$? We denote this contribution by $\phi_{i}(\tilde{\pi}_{s}^{a})$. In other words, how can we fairly distribute—across the state features—the difference between the action probability when all feature values are known and the action probability when none are known? Below, we define four axioms that any attribution $\phi_{i}(\tilde{\pi}_{s}^{a})$, for all $i\in\mathcal{F}$, should satisfy to qualify as a fair credit allocation. When referring to a feature’s contribution, we implicitly mean the contribution of its value $s_{i}$ at state $s$.
  • Axiom 1 (Efficiency). The difference in value between the action probability when all features are known, $\tilde{\pi}_{s}^{a}({\cal F})$, and the action probability when no features are known, $\tilde{\pi}_{s}^{a}(\emptyset)$, is fully distributed over the features ${\cal F}$:
    $\tilde{\pi}_{s}^{a}({\cal F})-\tilde{\pi}_{s}^{a}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{\pi}_{s}^{a})$
  • Axiom 2 (Linearity). For a composite policy $\alpha\,\pi+(1-\alpha)\,\omega$ that follows, at each decision stage, policy $\pi$ with probability $\alpha$ and policy $\omega$ with probability $(1-\alpha)$, a state feature’s contribution to the action probability $\alpha\,\pi(s,a)+(1-\alpha)\,\omega(s,a)$ is equal to the weighted sum of its contributions to the individual action probabilities $\pi(s,a)$ and $\omega(s,a)$:
    $\phi_{i}(\alpha\,\tilde{\pi}_{s}^{a}+(1-\alpha)\,\tilde{\omega}_{s}^{a})=\alpha\,\phi_{i}(\tilde{\pi}_{s}^{a})+(1-\alpha)\,\phi_{i}(\tilde{\omega}_{s}^{a}).$
  • Axiom 3 (Symmetry). Two features are assigned the same contribution to the action probability $\pi(s,a)$ if they make equal marginal contributions to the probability of selecting action $a$ in state $s$ for all possible coalitions of features:
    $\phi_{i}(\tilde{\pi}_{s}^{a})=\phi_{j}(\tilde{\pi}_{s}^{a})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{\pi}_{s}^{a}(\mathcal{C}\cup\{i\})=\tilde{\pi}_{s}^{a}(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}.$
  • Axiom 4 (Nullity). A feature is assigned zero contribution to the action probability $\pi(s,a)$ if it makes zero marginal contribution to the probability of selecting action $a$ in state $s$ for all possible coalitions of feature values:
    $\phi_{i}(\tilde{\pi}_{s}^{a})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{\pi}_{s}^{a}(\mathcal{C}\cup\{i\})=\tilde{\pi}_{s}^{a}(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}.$
These four axioms are directly analogous to the standard axioms from cooperative game theory used to fairly assign player contributions to the outcome of a coalitional game, as discussed in Section 2.
Proposition 2
The Shapley values $\phi_{1}(\tilde{\pi}_{s}^{a}),\dots,\phi_{n}(\tilde{\pi}_{s}^{a})$ for the discrete policy game $({\cal F},\tilde{\pi}_{s}^{a})$ are the unique allocation of the action probability $\pi(s,a)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The proofs for all propositions are presented in Appendix A. The Shapley values of the discrete policy game provide a principled way to attribute action probabilities to individual state features, offering a concrete explanation of agent behaviour.
We now revisit the example introduced in Sections 3 and 1, where an autonomous vehicle navigates through junctions using road signs that indicate both the direction and distance to the destination. At the first junction, the vehicle turns right. How can we explain this behaviour? We apply our framework to attribute the decision to the observed features at that state. However, we must first consider how the agent might behave when one or more features are unknown.
To explore this, we look at two possible ways the agent could act at the first junction when features are unknown. For clarity, we refer to these behaviours as separate agents. Agent A bases its decisions solely on the direction the sign indicates; if the direction is visible, it follows it. Otherwise, it chooses randomly. Agent B uses both direction and distance to determine which way to turn; if the direction or distance is visible in the first state, it moves right. Otherwise, it also chooses randomly. These assumptions give rise to different policies under unknown information, shown in Table 1.
Table 1: Agent behaviour under different subsets of observed features. Each cell shows the probability of turning left (L) and right (R).
Agent Both features Direction only Distance only No features
A $\pi(\text{L})=0,\ \pi(\text{R})=1$ 0, 1010,\ 10 , 1 0.5, $,\ 0.5$ 0.5 0.5, $,\ 0.5$ 0.5
B 0, 1010,\ 10 , 1 0, 1010,\ 10 , 1 0, 1010,\ 10 , 1 0.5, $,\ 0.5$ 0.5
We then compute the Shapley values for each policy, attributing the agent’s choice of action to the direction and distance features. The results are shown in Table 2. For Agent A, all credit is assigned to the direction feature since the policy disregards distance entirely. For Agent B, both features contribute equally because they both support the decision to turn right. In both cases, the features that increase the probability of turning right also decrease the probability of turning left—resulting in negative Shapley values. This example illustrates how the explanation depends critically on what is assumed about agent behaviour under partial information. We next discuss how we might approach this problem in general.
Table 2: Shapley values for each feature under the two agent policies.
Agent Action $\phi_{\text{direction}}$ $\phi_{\text{distance}}$
A L $-0.5$ 00
R $0.5=\tfrac{1}{2}[(1-0.5)+(1-0.5)]$ 00
B L $-0.25$ $-0.25$
R $0.25$ $0.25$
4.2 What Can Be Assumed About Agent Behaviour When Some State Features Are Unknown?
We now consider how an agent behaves when some state features are unknown—a question that must be addressed to compute Shapley values. The agent’s policy $\pi$ defines behaviour only when all feature values are known; it does not specify how the agent acts under partial information. However, to evaluate the characteristic function $\tilde{\pi}_{s}^{a}(\mathcal{C})$ for all subsets $\mathcal{C}\subseteq{\cal F}$, we must make assumptions about behaviour in these cases. In what follows, we introduce one such approach and then discuss alternative modelling choices and their implications.
Definition 3
Let $\mu:{\cal S}_{\mathcal{C}}\rightarrow\Delta({\cal A})$ be a policy defined over a subset of features $\mathcal{C}$ such that the probability of selecting action $a$ when the values of features $\bar{\mathcal{C}}$ are unknown is the expected probability of selecting $a$ given the known feature values $s_{\mathcal{C}}$:
$\mu(s_{\mathcal{C}},a)=\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\pi(s^{\prime},a),$
where $s_{\mathcal{C}}$ denotes the values of the features in $\mathcal{C}$ at state $s$. Here, ${\cal S}^{+}$ is the set of non-terminal states. The distribution $p^{\pi}(S\mid S_{\mathcal{C}}=s_{\mathcal{C}})$ is the conditional steady-state distribution: the probability that an agent following policy $\pi$ is in state $s$, conditioned on observing $s_{\mathcal{C}}$.
Among all policies defined over the feature subset $\mathcal{C}$, the policy $\mu$ is the one that deviates the least from $\pi$ in expectation when comparing action probabilities.
Proposition 4
When comparing policies using mean squared error between the probabilities of selecting action $a$, the policy $\mu$ is the policy that deviates the least in expectation from the policy $\pi$:
$\mu=\operatorname*{arg\,min}_{g}\mathbb{E}\left[\lVert\pi(S,a)-g(S_{\mathcal{C}},a)\rVert^{2}\right],$ (5)
where $g:{\cal S}_{\mathcal{C}}\rightarrow\Delta({\cal A})$ is also defined over the subset of features $\mathcal{C}$.
This motivates our definition of the characteristic function $\tilde{\pi}_{s}^{a}(\mathcal{C})$: the probability of selecting action $a$ at state $s$ when only the features in $\mathcal{C}$ are known is taken to be the value given by the policy $\mu$ at that state.
Definition 5 (Policy Characteristic Function)
The probability of selecting action $a$ when the values of features $\bar{\mathcal{C}}$ are unknown is defined as the expected probability of selecting $a$ given the known feature values $s_{\mathcal{C}}$:
$\tilde{\pi}_{s}^{a}(\mathcal{C})\equiv\mu(s_{\mathcal{C}},a)=\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\pi(s^{\prime},a).$
Remark 6
By substituting $\mathcal{C}=\emptyset$ into $\tilde{\pi}_{s}^{a}(\mathcal{C})$, the probability of choosing action $a$ when all features are unknown is given by:
$\tilde{\pi}_{s}^{a}(\emptyset)=\mathbb{E}[\pi(S,a)].$
The expectation is taken over the marginal steady-state distribution $p^{\pi}(S)$—the probability that an agent following policy $\pi$ is in state $s$.
When only the features in $\mathcal{C}$ are known, the policy $\mu$ provides the closest approximation to the original policy $\pi$. As such, the difference between $\mu(s_{\mathcal{C}},a)$ and $\pi(s,a)$ quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are to reproduce the behaviour of policy $\pi$; this reflects the influence of those unknown features. Defining the characteristic function as $\tilde{\pi}_{s}^{a}(\mathcal{C})=\mu(s_{\mathcal{C}},a)$ thus ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the policy game. Because this construction relies only on the agent’s policy, not on how it was derived, it is general to any policy—optimal and non-optimal.
4.3 How Else Could Agent Behaviour Be Represented Under Partial Information?
In the previous subsection, we defined behaviour with unknown features by marginalising the policy over the conditional steady-state distribution. Here, we propose and briefly discuss several alternative approaches.
One approach is to learn policies for each subset of observed features. However, policies learnt with partial information may fail to converge or meaningfully reflect the original policy we aim to explain. Another approach is to treat “unknown” as a valid feature value, but this would require modifying the MDP so the agent can learn to interpret the “unknown” feature appropriately. Similarly, unknown features could be imputed with a placeholder, but without adapting training, the agent has no reason to interpret such values as indicating missingness. These approaches all require changes to the MDP or training process, making them restrictive and unsuitable for post hoc explanation—though they may offer promising directions for future work.
Another option is to define the characteristic function by marginalising over the joint distribution of feature values. Unlike the approaches above, this does not require changing the MDP or training process. Formally, it is defined as:
$\tilde{\pi}_{s}^{a}(\mathcal{C})=\mathbb{E}[\pi(s_{\mathcal{C}},S_{\bar{\mathcal{C}}},a)]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime})\pi(\tau(s,s^{\prime},\mathcal{C}),a),$ (6)
where $\tau(s,s^{\prime},\mathcal{C})_{i}=s_{i}$ if $i\in\mathcal{C}$, and $s^{\prime}_{i}$ otherwise. Unknown feature values are replaced with values sampled from the joint marginal distribution over states.
This approach differs from marginalising over the conditional steady-state distribution in much the same way as in supervised learning (Section 2.3). It assumes feature independence and yields equivalent results when that assumption holds. As a result, it assigns zero Shapley value to any feature not explicitly used by the policy. This outcome may appear desirable, but it is less meaningful in many machine learning models, where all features and interdependencies are explicitly used. When features are dependent, joint marginalisation may produce combinations of values that do not correspond to any valid state. Since Shapley value attribution relies on measuring behavioural changes in response to changes in known feature values, such undefined policy queries make these changes uninterpretable. For this reason, we focus on the conditional approach in this work, although adapting our framework for joint marginalisation is straightforward. The choice of characteristic function may ultimately depend on the user, who must understand the underlying assumptions to interpret the resulting Shapley values correctly.
4.4 An Illustrative Example
[…]
We now apply the proposed framework to explain the agent’s behaviour in these two states, representing behaviour under unknown features by marginalising over the conditional steady-state distribution, as defined in Section 4.2. To do so, we first compute the steady-state distribution of the policy. Starting in State 1, the agent turns right, then left at State 2, and terminates thereafter. This yields a steady-state distribution of $0.5$ for each of the two states.
Using this distribution, we compute the agent’s behaviour under all subsets of observed features and use the resulting characteristic values to calculate Shapley values for each action. These are shown in Table 3.
Table 3: Characteristic values and Shapley values for each state-action pair $(s,a)$, computed by marginalising over the conditional steady-state distribution. Each state is denoted by its arrow direction (R for right, L for left) and the distance to the goal.
State Action $p^{\pi}(s)$ $\tilde{\pi}_{s}^{a}(\text{Both})$ $\tilde{\pi}_{s}^{a}(\text{Dir})$ $\tilde{\pi}_{s}^{a}(\text{Dist})$ $\tilde{\pi}_{s}^{a}(\emptyset)$ $\phi_{\text{Dir}}$ $\phi_{\text{Dist}}$
$(\text{R},10)$ L $0.5$ 00 00 00 $0.5$ $-0.25$ $-0.25$
[…]
R 00 00 00 $0.5$ $-0.25$ $-0.25$
The characteristic values and Shapley values are identical to those of Agent B in the earlier example: direction and distance contribute equally to the action taken in both visited states. This reflects our assumption—implicit in marginalising over the conditional steady-state distribution—that the agent uses all available feature information when acting. Accordingly, both features are treated as necessary for reproducing the agent’s policy, and responsibility is attributed proportionally.
This example demonstrates how defining the characteristic function by marginalising over the conditional steady-state distribution enables Shapley values to provide meaningful explanations of agent behaviour. [^2] See Section B.1 for an additional example illustrating how SVERL explains agent behaviour.
4.5 Explaining Behaviour in Continuous Action Spaces
So far, we have considered agent behaviour in terms of how feature values influence the probabilities of selecting discrete actions. In continuous action spaces, however, the probability of selecting any particular action is generally zero. As such, policies cannot be interpreted as assigning probabilities to individual actions, and the approach developed in the previous subsections does not directly apply. Instead, we consider an alternative perspective tailored to continuous action settings.
A common policy representation in continuous control is a fixed-variance Gaussian, where actions are sampled from a distribution $\mathcal{N}(\mu(s),\sigma^{2})$, where $\mu:{\cal S}\rightarrow\mathbb{R}$ maps each state to the mean of the distribution. In this setting, a sound approach to explaining agent behaviour is to analyse how feature values determine the policy’s expected action $\mu(s)$. For example, how does a pendulum’s angular velocity influence the expected torque an agent applies in a pendulum balancing task?
This approach naturally leads to the question: how much does each individual feature contribute to the expected action $\mu(s)$? As before, we take a game-theoretic perspective, viewing $\mu(s)$ as the outcome of a coalitional game played by the state features. We identify Shapley values as the unique solution that satisfies four axioms defining a fair credit allocation.
Definition 7 (Continuous Policy Game)
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\tilde{\mu}_{s}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\tilde{\mu}_{s}(\mathcal{C})$ returns the expected action in state $s$ when only the features in $\mathcal{C}\subseteq{\cal F}$ are known.
Central to the continuous policy game $({\cal F},\tilde{\mu}_{s})$ is the characteristic function $\tilde{\mu}_{s}$, which extends the mean function $\mu$ to situations where only a subset of state features is known. The precise construction of $\tilde{\mu}_{s}$ will be addressed later in this section; for now, we assume it is given and turn to our central question: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the policy’s expected action $\mu(s)$? We denote this contribution by $\phi_{i}(\tilde{\mu}_{s})$. In other words, how can we fairly allocate—across the state features—the difference between the expected action when all features are known and when none are? Below, we define four more axioms that any attribution $\phi_{i}(\tilde{\mu}_{s})$, for all $i\in\mathcal{F}$, should satisfy to qualify as a fair credit allocation.
  • Axiom 1 (Efficiency). The difference in value between the expected action when all features are known, $\tilde{\mu}_{s}({\cal F})$, and the expected action when no features are known, $\tilde{\mu}_{s}(\emptyset)$, is fully distributed across the features:
    $\tilde{\mu}_{s}({\cal F})-\tilde{\mu}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{\mu}_{s})$
  • Axiom 2 (Linearity). For a composite policy that samples actions from a Gaussian whose mean $\alpha\,\tilde{\mu}(s)+(1-\alpha)\,\tilde{\nu}(s)$ is the weighted average of the mean functions $\tilde{\mu}(s)$ and $\tilde{\nu}(s)$, a feature’s contribution to the combined policy’s expected action is equal to the weighted sum of its contributions to each individual policy’s expected action:
    $\phi_{i}(\alpha\,\tilde{\mu}_{s}+(1-\alpha)\,\tilde{\nu}_{s})=\alpha\,\phi_{i}(\tilde{\mu}_{s})+(1-\alpha)\,\phi_{i}(\tilde{\nu}_{s})$
  • Axiom 3 (Symmetry). Two features are assigned the same contribution to the expected action if they make equal marginal contributions for all possible coalitions:
    $\phi_{i}(\tilde{\mu}_{s})=\phi_{j}(\tilde{\mu}_{s})\quad\text{if}\quad\tilde{\mu}_{s}(\mathcal{C}\cup\{i\})=\tilde{\mu}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature is assigned zero contribution if it makes zero marginal contribution to the expected action for all possible coalitions:
    $\phi_{i}(\tilde{\mu}_{s})=0\quad\text{if}\quad\tilde{\mu}_{s}(\mathcal{C}\cup\{i\})=\tilde{\mu}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
These four axioms are directly analogous to those used in the discrete case and identify the Shapley values as the unique solution for attributing the mean $\mu(s)$ across features.
Proposition 8
The Shapley values $\phi_{1}(\tilde{\mu}_{s}),\dots,\phi_{n}(\tilde{\mu}_{s})$ for the continuous policy game $({\cal F},\tilde{\mu}_{s})$ are the unique allocation of the expected action $\mu(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the continuous policy game provide a principled way to attribute an agent’s expected action to individual state features, offering a concrete explanation of behaviour in continuous action spaces. As in the discrete case, computing Shapley values requires defining the characteristic function $\tilde{\mu}_{s}(\mathcal{C})$, which entails making assumptions about how the agent behaves when only partial state information is available.
Definition 9
Let $\zeta:{\cal S}_{\mathcal{C}}\rightarrow\Delta({\cal A})$ be a policy defined over a subset of features $\mathcal{C}$, where actions are sampled from a fixed-variance Gaussian $\mathcal{N}(\omega(s_{\mathcal{C}}),\sigma^{2})$. The mean function $\omega:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}$ defines the expected action when only the features in $\mathcal{C}$ are known:
$\omega(s_{\mathcal{C}})=\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\,\mu(s^{\prime}).$
Among all fixed-variance Gaussian policies defined over $\mathcal{C}$, the policy $\zeta$ is the one whose mean $\omega$ deviates the least from $\mu$ in expectation.
Proposition 10
When comparing mean functions using mean squared error, the function $\omega$ deviates the least in expectation from the mean $\mu$ of the full policy $\pi$:
$\omega=\operatorname*{arg\,min}_{g:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}}\mathbb{E}\left[\lVert\mu(S)-g(S_{\mathcal{C}})\rVert^{2}\right].$ (7)
This motivates our definition of the characteristic function $\tilde{\mu}_{s}(\mathcal{C})$: the expected action when only the features in $\mathcal{C}$ are known is taken to be the value given by $\omega(s_{\mathcal{C}})$ at that state.
Definition 11 (Continuous Policy Characteristic Function)
The expected action when only the features in $\mathcal{C}$ are known is defined as:
$\tilde{\mu}_{s}(\mathcal{C})\equiv\omega(s_{\mathcal{C}})=\mathbb{E}[\mu(S)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\,\mu(s^{\prime}).$ (8)
When only the features in $\mathcal{C}$ are known, the function $\tilde{\mu}_{s}(\mathcal{C})$ provides the closest approximation to the full policy’s expected action $\mu(s)$. As such, the difference between $\tilde{\mu}_{s}(\mathcal{C})$ and $\mu(s)$ quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are to reproduce the agent’s behaviour. Defining the characteristic function this way ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the continuous policy game. This construction is general to any policy.
As in the discrete setting, alternative characteristic functions may also be considered here. In particular, a marginal variant of $\tilde{\mu}_{s}(\mathcal{C})$—which assumes feature independence—could be used in place of the conditional formulation discussed above.
We do not present an illustrative example in this setting, as doing so would require estimating the steady-state distribution and conditional expectations in continuous state-spaces—a task beyond this work’s scope. Nonetheless, the theory developed here offers a general and principled foundation for future approaches to explaining agent behaviour in continuous action domains.
5 Explaining Performance
We now turn to explaining performance. In reinforcement learning, the performance of a policy $\pi$ is quantified by the state-value function $v^{\pi}:{\cal S}\rightarrow\mathbb{R}$, which gives the expected return from each state when following that policy:
$v^{\pi}(s)\coloneqq\mathbb{E}_{\pi}\left[G_{t}\mid S_{t}=s\right].$ (9)
Given that states are composed of feature values, it is natural to ask how individual features contribute to this expected return. For example, how does a navigation agent’s $x-$coordinate influence the expected return of its policy at a given location? In this section, we develop a theoretical framework for attributing expected return to the values of state features, offering a structured explanation of agent performance under policy $\pi$.
5.1 State Features Collaborate to Determine Expected Return
As with explaining agent behaviour, we start by taking a game-theoretic view of expected return. Just as an agent’s policy arises from the combined influence of its state features, the expected return $v^{\pi}(s)$ can also be viewed as the outcome of a collaborative process among those features. That is, the return achieved from a state is affected by how the policy responds to the information available at that state. This perspective motivates the question: how much does each individual feature contribute to the overall return? To answer this, we define another coalitional game over state features, introduce a new set of fairness axioms, and show that Shapley values offer a unique principled solution.
Definition 12 (Performance Game)
A set =${\cal F}=\{1,\dots,n\}$ of state features and a characteristic function $\tilde{v}^{\pi}_{s}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\tilde{v}^{\pi}_{s}(\mathcal{C})$ denotes the expected return from state $s$ when policy $\pi$ has access to only the values of features in $\mathcal{C}\subseteq{\cal F}$.
The characteristic function $\tilde{v}^{\pi}_{s}$ extends the value function $v^{\pi}$ to settings where only a subset of feature values are accessible. As before, we postpone the precise construction of $\tilde{v}^{\pi}_{s}(\mathcal{C})$ until later. For now, we ask: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the expected return $v^{\pi}(s)$? We denote this by $\phi_{i}(\tilde{v}^{\pi}_{s})$, define a set of fairness axioms that any such attribution should satisfy, and identify the Shapley values as the unique solution.
  • Axiom 1 (Efficiency). The difference in expected return between the policy with all feature values known and the policy with none known is fully distributed across the features:
    $\tilde{v}^{\pi}_{s}({\cal F})-\tilde{v}^{\pi}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{v}^{\pi}_{s})$
  • Axiom 2 (Linearity). For a composite policy $\alpha\,\pi+(1-\alpha)\,\omega$ that follows $\pi$ with probability $\alpha$ and $\omega$ with probability $1-\alpha$, a feature’s contribution to the expected return of the composite policy is equal to the same weighted combination of its contributions under each component policy:
    $\phi_{i}(\alpha\,\tilde{v}^{\pi}_{s}+(1-\alpha)\,\tilde{v}^{\omega}_{s})=\alpha\,\phi_{i}(\tilde{v}^{\pi}_{s})+(1-\alpha)\,\phi_{i}(\tilde{v}^{\omega}_{s})$
  • Axiom 3 (Symmetry). Two features are assigned the same contribution if they make equal marginal contributions to the expected return for all possible coalitions:
    $\phi_{i}(\tilde{v}^{\pi}_{s})=\phi_{j}(\tilde{v}^{\pi}_{s})\quad\text{if}\quad\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature is assigned zero contribution if it makes no marginal contribution to the expected return for all possible coalitions:
    $\phi_{i}(\tilde{v}^{\pi}_{s})=0\quad\text{if}\quad\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\tilde{v}^{\pi}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
These four axioms mirror those used to explain behaviour and are directly analogous to the standard axioms from cooperative game theory.
Proposition 13
The Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ for the performance game $({\cal F},\tilde{v}^{\pi}_{s})$ are the unique allocation of the expected return $v^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the performance game provide a principled way to attribute expected return to individual state features, offering a concrete explanation of agent performance at a given state.
[…]
We apply our framework to attribute the expected return at each state to its observed feature values. However, we must first make assumptions about how performance is affected by unknown features. We assume this corresponds to acting with partial information in the given state (at any decision step) and full observability elsewhere.
To specify how the agent acts with partial information, we reuse the two agents introduced earlier: Agent A follows the direction if visible and acts randomly otherwise. Agent B uses both direction and distance when available and acts randomly when neither is present. These assumptions define the policy under partial information and thus determine the expected return from each state for every feature subset. The resulting characteristic and Shapley values are shown in Tables 4 and 5.
Table 4: Characteristic values and Shapley values for expected return under Agent A.
State $(s)$ $\tilde{v}^{\pi}_{s}(\text{Both})$ $\tilde{v}^{\pi}_{s}(\text{Dir})$ $\tilde{v}^{\pi}_{s}(\text{Dist})$ $\tilde{v}^{\pi}_{s}(\emptyset)$ $\phi_{\text{Dir}}$ $\phi_{\text{Dist}}$
$(\text{R},10)$ $8$ $8$ $8$ $8$ 00 00
$(\text{L},2)$ $9$ $9$ $4$ $4$ $5$ 00
Table 5: Characteristic values and Shapley values for expected return under Agent B.
State $(s)$ $\tilde{v}^{\pi}_{s}(\text{Both})$ $\tilde{v}^{\pi}_{s}(\text{Dir})$ $\tilde{v}^{\pi}_{s}(\text{Dist})$ $\tilde{v}^{\pi}_{s}(\emptyset)$ $\phi_{\text{Dir}}$ $\phi_{\text{Dist}}$
$(\text{R},10)$ $8$ $8$ $8$ $8$ 00 00
$(\text{L},2)$ $9$ $9$ $9$ $4$ $2.5$ $2.5$
Consider first the state $(\text{R},10)$. For both agents, neither feature contributes to performance. The reason is that turning left or right from the initial junction results in paths of equal length to the goal, yielding identical returns. In contrast, in our explanation of behaviour, the direction feature received credit for prompting the agent to turn right. This example demonstrates a key distinction between explaining behaviour and performance: a feature can influence behaviour without affecting performance.
In the second state, $(\text{L},2)$, the features do influence performance. For Agent A, all credit goes to the direction feature, reflecting its exclusive influence on the agent’s decision. For Agent B, both features contribute equally because either can be used to determine the correct action. These performance attributions mirror our behaviour analysis under the same assumptions, illustrating how the two forms of explanation can often align.
This example demonstrates how the proposed framework explains performance by attributing expected returns to individual state features. Crucially, the resulting explanations depend on the agent’s assumed behaviour under partial information. The following section outlines a general approach to addressing this challenge.
5.2 What Can Be Assumed About Performance When Some State Features Are Unknown?
We now consider how to define the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$: the expected return from state $s$ when the agent’s policy has access only to the features in $\mathcal{C}$. Since the value function $v^{\pi}(s)$ assumes full observability, it does not specify this quantity, so we must make additional assumptions. If the agent’s expected return from state $s$ changes when certain feature values are unknown, this must reflect a shift in the agent’s behaviour at that state, assuming all else remains constant. Motivated by this, we define performance under unknown information as the expected return when the agent acts with partial information in state $s$ and follows the fully observed policy elsewhere. This definition isolates the contribution of features at state $s$ and allows us to express $\tilde{v}^{\pi}_{s}(\mathcal{C})$ using the policy characteristic function $\tilde{\pi}_{s}^{a}(\mathcal{C})$ introduced in Section 4.2.
Recall that $\tilde{\pi}_{s}^{a}(\mathcal{C})$ denotes the probability of selecting action $a$ in state $s$ when only the features in $\mathcal{C}$ are known:
$\tilde{\pi}_{s}^{a}(\mathcal{C})=\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}=s_{\mathcal{C}}].$
These probabilities form a complete action distribution at $s$ under partial information, which we can use to evaluate performance with unknown features.
Definition 14 (Performance Characteristic Function)
The expected return from state $s$ when an agent has access only to features in $\mathcal{C}$ is defined as:
$\tilde{v}^{\pi}_{s}(\mathcal{C})=\mathbb{E}_{\mu}\left[G_{t}\mid S_{t}=s\right],$ (10)
where the policy $\mu$ is defined as
$\mu(s_{t},a_{t})=\begin{cases}\pi_{s_{t},a_{t}}(\mathcal{C})&\text{if }s_{t}=s,\\ \pi(s_{t},a_{t})&\text{otherwise.}\end{cases}$
This definition ensures that unknown features modify only behaviour at state $s$, isolating their contribution to performance. The agent behaves according to its original policy $\pi$ in all other states.
Remark 15
Since we defined $\tilde{\pi}_{s}^{a}(\mathcal{C})$ as a marginal over the full state space, it may assign nonzero probability to unavailable actions in $s$. The distribution over ${\cal A}(s)$ must be renormalised in such cases.
When only the features in $\mathcal{C}$ are available, the expected return $\tilde{v}^{\pi}_{s}(\mathcal{C})$ captures the outcome of following the best approximation to the original policy $\pi$ under this partial information. The difference between $\tilde{v}^{\pi}_{s}(\mathcal{C})$ and $v^{\pi}(s)$ thus quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are for achieving the original expected return. Defining the characteristic function in this way ensures that the differences across feature subsets used to compute Shapley values capture the influence of the unknown features, making it a principled choice for the performance game. As before, this construction is general to any policy.
Alternative definitions of agent behaviour under partial information could also be substituted into our proposed characteristic function. For example, the marginal variant of $\tilde{\pi}_{s}^{a}(\mathcal{C})$—which assumes feature independence—may be appropriate in some settings; see Section 4.3 for further discussion. Similarly, in continuous action spaces, the policy with partial information used in performance estimation can be defined as a Gaussian with mean $\mu_{s}(\mathcal{C})$, in place of the discrete policy based on $\tilde{\pi}_{s}^{a}(\mathcal{C})$; this substitution naturally extends the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$ to continuous domains. We leave a detailed treatment of this extension to future work.
We do not present a new illustrative example for the proposed characteristic function because its outcome is identical to the example in the previous subsection. In the road-sign domain, we noted how the policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$ behaves exactly like Agent B under partial information. Since the performance characteristic $\tilde{v}^{\pi}_{s}(\mathcal{C})$ is defined using this same policy characteristic, it leads to the same values and attributions already shown for Agent B in Table 5. As such, the example already demonstrates how Shapley values enable faithful explanations of agent performance using this characteristic function. We provide an additional example illustrating how SVERL explains agent performance in Section B.2.
6 Explaining Value Estimation
We now turn to our third and final element of reinforcement learning that we will explain: value estimation. This element refers to how well an agent, or an observer, can estimate $v^{\pi}(s)$: the expected return of a policy $\pi$ from a given state. As before, our goal is to understand how the values of individual features contribute to this estimate.
At first glance, value estimation may seem indistinguishable from explaining performance: both involve analysing the expected return from a state. However, there is a crucial difference. When explaining performance, we asked how feature values influence the return an agent receives—by potentially altering the agent’s behaviour when some features are unknown. When explaining value estimation, by contrast, we ask how feature values influence the return an agent is estimated to receive—without altering how the agent behaves in that state.
To make this distinction precise, we introduce $\hat{v}^{\pi}:{\cal S}\rightarrow\mathbb{R}$ to denote expected return estimates from each state under policy $\pi$. When the estimate is accurate, $\hat{v}^{\pi}(s)$ equals $v^{\pi}(s)$, the true expected return from $s$. By changing our framing to treat $\hat{v}^{\pi}$ as a predictive function rather than an outcome, we present a theoretical analysis of explaining value estimation using the influence of feature values.
6.1 State Features Collaborate to Determine Value Estimates
We now take a game-theoretic view of value estimation. The estimated value $\hat{v}^{\pi}(s)$ can also be viewed as the outcome of a collaborative process among features. That is, each feature value in $s$ contributes to the estimation of return under policy $\pi$. To attribute this estimate across the feature values, we define another coalitional game over the features, introduce a new set of fairness axioms tailored to the estimation setting, and identify the Shapley values as the unique solution.
Definition 16 (Value Estimation Game)
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\hat{v}^{\pi}_{s}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\hat{v}^{\pi}_{s}(\mathcal{C})$ denotes the estimated expected return when only the feature values $s_{\mathcal{C}}\in{\cal S}_{\mathcal{C}}$ are observed.
As before, we postpone the precise construction of the characteristic function $\hat{v}^{\pi}_{s}$ until later. For now, we ask: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the estimate $\hat{v}^{\pi}(s)$? We denote this contribution by $\phi_{i}(\hat{v}^{\pi}_{s})$, and define four axioms that any such attribution should satisfy.
  • Axiom 1 (Efficiency). The difference in value between the estimated return when all features are known, $\hat{v}^{\pi}_{s}({\cal F})$, and when none are known, $\hat{v}^{\pi}_{s}(\emptyset)$, is fully distributed across the features:
    $\hat{v}^{\pi}_{s}({\cal F})-\hat{v}^{\pi}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\hat{v}^{\pi}_{s})$
  • Axiom 2 (Linearity). For a composite policy $\alpha\,\pi_{1}+(1-\alpha)\,\pi_{2}$, a feature’s contribution to the estimated expected return under the combined policy is equal to the same weighted sum of its contributions under each individual policy:
    $\phi_{i}(\alpha V^{\pi_{1}}_{s}+(1-\alpha)V^{\pi_{2}}_{s})=\alpha\,\phi_{i}(V^{\pi_{1}}_{s})+(1-\alpha)\,\phi_{i}(V^{\pi_{2}}_{s})$
  • Axiom 3 (Symmetry). Two features receive equal attribution if they make identical marginal contributions to the estimated return for all possible coalitions:
    $\phi_{i}(\hat{v}^{\pi}_{s})=\phi_{j}(\hat{v}^{\pi}_{s})\quad\text{if}\quad\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature receives zero attribution if it makes no marginal contribution to the estimated return in any coalition:
    $\phi_{i}(\hat{v}^{\pi}_{s})=0\quad\text{if}\quad\hat{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\hat{v}^{\pi}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
These four axioms mirror those used when explaining behaviour and performance.
Proposition 17
The Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ for the value estimation game $({\cal F},\hat{v}^{\pi}_{s})$ are the unique allocation of the estimated return $\hat{v}^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the value estimation game provide a principled way to attribute state-value estimates to individual state features, offering a structured explanation of value estimation. As with the previous coalitional games, computing these Shapley values requires precisely defining the characteristic function $\hat{v}^{\pi}_{s}(\mathcal{C})$, which we turn to next.
6.2 What Can Be Assumed About Value Estimates When Some State Features Are Unknown?
We now consider how to define the characteristic function $\hat{v}^{\pi}_{s}(\mathcal{C})$: the estimated return from state $s$ when only the features in $\mathcal{C}\subseteq{\cal F}$ are observed. A natural principle is to estimate value using only the information that is actually available. The estimated return should reflect the best assessment based on the observed features alone.
Definition 18
Let $\hat{u}^{\pi}:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}$ be a value function defined over a subset of features $\mathcal{C}\subseteq{\cal F}$ such that the estimated return from state $s$ when only the features in $\mathcal{C}$ are observed is given by:
$\hat{u}^{\pi}(s_{\mathcal{C}})=\mathbb{E}[\hat{v}^{\pi}(S)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\,\hat{v}^{\pi}(s^{\prime}).$
Among all functions defined over the observed features, $\hat{u}^{\pi}$ is the one that deviates the least, in expectation, from the full-information estimate $\hat{v}^{\pi}$.
Proposition 19
When comparing return estimates using mean squared error, the function $\hat{u}^{\pi}$ is the closest in expectation to $\hat{v}^{\pi}$ among all functions defined over the subset of features $\mathcal{C}$:
$\hat{u}^{\pi}=\operatorname*{arg\,min}_{g:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}}\mathbb{E}\left[\lVert\hat{v}^{\pi}(S)-g(S_{\mathcal{C}})\rVert^{2}\right].$
This motivates our definition of the characteristic function for value estimation:
Definition 20 (Value Estimation Characteristic Function)
The estimated return from state $s$ when only the features in $\mathcal{C}$ are observed is given by:
$\hat{v}^{\pi}_{s}(\mathcal{C})\coloneqq\hat{u}^{\pi}(s_{\mathcal{C}})=\mathbb{E}[\hat{v}^{\pi}(S)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\,\hat{v}^{\pi}(s^{\prime}).$ (11)
When only the features in $\mathcal{C}$ are observed, the function $\hat{u}^{\pi}$ provides the best approximation of the full-information value estimate $\hat{v}^{\pi}(s)$. As such, the difference between $\hat{u}^{\pi}(s_{\mathcal{C}})$ and $\hat{v}^{\pi}(s)$ quantifies how necessary the unobserved features $\bar{\mathcal{C}}$ are for recovering the original estimate. Defining the characteristic function as $\hat{v}^{\pi}_{s}(\mathcal{C})=\hat{u}^{\pi}(s_{\mathcal{C}})$ ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the value estimation game.
As with explaining policy and performance, this construction is general to any policy. Alternative characteristic functions could also be used. In particular, a marginal variant of $\hat{v}^{\pi}_{s}(\mathcal{C})$—which assumes feature independence—may be suitable in some settings. An analogous discussion of this alternative appears for the policy characteristic function in Section 4.3.
6.3 Illustrative Example
[…]
  • $(\text{R},10)$: The initial state with a right-facing sign and distance 10 miles.
  • $(\text{L},2)$: The second state, reached by turning right, with a left-facing sign and distance 2 miles.
Using the characteristic function introduced above, we now apply our framework to analyse how value is estimated in each state. Table 6 presents the return estimates for different feature subsets, and the resulting Shapley values for each feature.
Table 6: The steady-state distribution, characteristic values, and Shapley values when explaining value estimation for each state in the road sign example (Figure 1).
State ($s$) $p^{\pi}(s)$ $\hat{v}^{\pi}_{s}(\text{Both})$ $\hat{v}^{\pi}_{s}(\text{Dir})$ $\hat{v}^{\pi}_{s}(\text{Dist})$ $\hat{v}^{\pi}_{s}(\emptyset)$ $\phi_{\text{Dir}}$ $\phi_{\text{Dist}}$
$(\text{R},10)$ $0.5$ $8$ $8$ $8$ $8.5$ $-0.25$ $-0.25$
[…]
We start by focusing on the contribution of the Distance feature. In state $(\text{R},10)$, the agent is 10 miles from the goal. The distance feature reduces the return estimate relative to the baseline with no features, resulting in a negative Shapley value ($-0.25$). In state $(\text{L},2)$, the agent is much closer to the goal, and the distance feature increases the return estimate, producing a positive Shapley value ($+0.25$).
These attributions reflect an intuitive interpretation: when the agent is far from the goal, the expected return is lower; when it is near, the expected return is higher. The Shapley values capture how each feature contributes to the estimate, correctly assigning responsibility to the Distance feature in line with how it influences return estimation.
The Direction feature receives identical attributions because it also helps differentiate the two visited states. In this domain, the sign’s direction indirectly conveys distance: the policy only encounters the right-facing sign when far from the goal and the left-facing sign when close.
This example demonstrates how our framework enables Shapley values to provide meaningful explanations of value estimation, revealing how individual features guide our estimate of the agent’s expected outcomes. [^3] See Section B.3 for an additional example illustrating how SVERL explains value estimation.
7 Understanding SVERL Through Examples
We present two examples using SVERL to explain agent-environment interactions in practice. The first example shows how explanations of behaviour, performance, and value estimation can align or diverge to reveal complementary insights. The second example applies the framework in a larger domain, demonstrating its scalability to the feasible limits of exact Shapley value computation. Unless stated otherwise, all examples are undiscounted and explain optimal policies computed via Value Iteration (Sutton and Barto, 2018). Steady-state distributions are either computed exactly or accurately estimated from trajectories; expected returns for the performance characteristic are learnt using Value Iteration. [^4] Code is available at https://github.com/djeb20/sverl.
7.1 Mastermind
Figure 2: Shapley value attributions for three successive states in Mastermind (top to bottom). First column: episode progression. Second column: feature contributions to the probability of selecting the optimal action taken (behaviour). Third column: contributions to the agent’s performance. Fourth column: contributions to the agent’s return estimate (value estimation). Shapley values are scaled to lie between $-1$ and $1$.
The first domain we consider is Mastermind, based on the eponymous code-breaking board game. In Mastermind, at the start of each episode, the environment randomly samples one of four codes: $\mathrm{AA}$, $\mathrm{AB}$, $\mathrm{BA}$, or $\mathrm{BB}$, hidden from the agent. At each decision stage, the agent presents one of the codes. The environment then returns two types of feedback to indicate how well the agent’s code matches the hidden code:
  • Position clue: the number of letters that are both correct and in the correct position.
  • Misplaced clue: the number of correct letters that are in the wrong position.
These two clues are calculated independently, and a letter contributing to the position clue cannot also contribute to the misplaced clue. Incorrect codes incur a reward of $-1$ and correct codes incur a reward of 00. The game terminates at the end of three decision stages, unless the agent identifies the code correctly in an earlier decision stage. Each state is represented as a game board, with up to three agent codes and each corresponding pair of clues. Each agent code has four associated features: two for the code and two for the feedback it received. If fewer than three codes have been presented, the remaining features take on an empty value.
We use SVERL to analyse an example episode where the hidden code is $\mathrm{AB}$. Figure 2 presents the trajectory for the agent’s policy: deterministically selecting $\mathrm{AA}$, $\mathrm{BA}$, then $\mathrm{AB}$, and explanations of behaviour, performance, and value estimation.
We start with the initial decision stage, where the agent has not yet presented a code. In other words, the board is clear. The agent’s policy deterministically selects the action $\mathrm{AA}$. The explanation of behaviour highlights the empty bottom row—corresponding to the agent’s first code—as the largest contributor towards the probability of this action. This explanation reflects that the agent’s policy chooses $\mathrm{AA}$ precisely because it’s in the initial state where no prior information is available. Similarly, the explanation of value estimation highlights this row too, but for decreasing the agent’s estimated return. Again, this reflects that the empty bottom row indicates the environment is in the first state, with the lowest expected return. In contrast, the performance explanation attributes no responsibility to any feature, correctly indicating that all possible initial agent codes result in maximum expected return.
In the second state, the first code $\mathrm{AA}$ and its clues have revealed: one of the $\mathrm{A}$s is in the correct position, and the other should be a $\mathrm{B}$. The agent now selects $\mathrm{BA}$. The behaviour and performance explanations identify the position clue as the primary driver of action selection and return. While the misplaced clue is informative, it is 00 for the initial code $\mathrm{AA}$ with any code, and thus does not meaningfully influence the agent’s decisions. For value estimation, the position clue of $1$ receives a negative attribution: a value of 00 would have revealed the hidden code and yielded a higher expected return, so the observed value of $1$ lowers the estimate.
In the final state, the agent correctly proposes $\mathrm{AB}$. The previous code and clues uniquely determine the hidden code and the correct action. The features corresponding to the code and clues receive high attribution across all three explanations: they reveal the optimal move (behaviour), justify the return received (performance), and identify a state with maximal return (value estimation).
This example illustrates how SVERL captures the evolving role of information in decision making, outcomes, and predictions as the game progresses.
7.2 Minesweeper
Figure 3: Shapley value attributions for two successive states in Minesweeper (top to bottom). First column: episode progression. Second column: feature contributions to the probability of selecting the optimal action taken (behaviour). Third column: contributions to the agent’s performance. Fourth column: contributions to the agent’s return estimate (value estimation). Mines are hidden from the agent; markers “M1” and “M2” are shown for reference. Shapley values are scaled to lie between $-1$ and $1$.
Minesweeper is played on a $4\times 4$ grid with two hidden mines randomly placed at the start of each episode and all squares unopened. The agent selects unopened squares to reveal either a mine-free square with a number indicating how many mines total are in adjacent squares, or a mine, which ends the episode with a reward of $-1$. If an opened square shows 00, adjacent squares are opened recursively until reaching nonzero borders. If all safe squares are revealed, the episode ends with a reward of 00. This domain presents a more complex and higher-dimensional challenge. Each state is represented as a game board, consisting of 16 features, each taking one of the following values: the number of adjacent mines (0, 1, or 2) or the value unopened. The domain has approximately $175,000$ states. We explain the behaviour, performance, and value estimation of a policy learnt via Q-learning (Watkins and Dayan, 1992), with the performance characteristic estimated using Monte Carlo rollouts. We consider the two states shown in Figure 3.
In the first state, the optimal action is to open square $(4,2)$. Many features provide information about the first mine (M1), but only the clue under square $(4,2)$ identifies the location of the second mine (M2). The clue is thus pivotal to safe and optimal play. The behaviour and performance explanations both assign high attribution to this feature. In Minesweeper, states with more unopened squares carry more uncertainty about mine locations and, therefore, have lower expected returns. This is reflected by the explanation of value estimation, where unopened squares receive negative attributions, and opened squares receive positive attributions.
In the second state, the agent opens square $(4,4)$. The clue revealed at $(4,2)$ by the previous action has uniquely identified both mine locations, determining the optimal action and return. The feature corresponding to the clue again receives high attribution in the behaviour and performance explanations. As before, opened squares contribute positively and unopened squares negatively to value estimation, reflecting how greater certainty leads to higher return estimates.
This example shows how our framework scales to more complex environments, revealing which features drive safe decision-making and how uncertainty affects expected outcomes.
8 Understanding and Interpreting SVERL
SVERL provides a lens through which we can understand how individual feature values contribute to an agent’s behaviour, performance, and value estimation. These insights go beyond what the policy or expected return reveals alone. While many of the conclusions drawn in our examples may seem intuitive in hindsight, they were not obvious when observing only the agent’s actions or value function. Even in well-understood domains, SVERL revealed how specific features shaped decision-making and outcomes, offering a more detailed understanding of the reinforcement learning problem.
In this section, we reflect on the insights provided by SVERL, considering how the three explanation types compare and contrast to form a richer picture of agent behaviour. We also discuss how these explanations should be interpreted in practice, including common pitfalls and the risk of misunderstanding.
8.1 Comparing Explanations
Each explanation in the proposed framework captures a distinct aspect of reinforcement learning. Explanations of behaviour reveal how features influence the agent’s action choices. Explanations of performance reveal how features influence expected return. Explanations of value estimation reveal how features influence the estimation of the expected return. Understanding how these perspectives agree—and when they diverge—is important for forming a general understanding of agent-environment interactions.
The final state of Mastermind shows an example of when explanations of behaviour and performance align. The same features that drive the agent’s decision to guess AB also determine the outcome: they reveal the correct code and lead to the highest possible return. This agreement may not be surprising, since performance is determined by behaviour. However, SVERL also demonstrates that this alignment between behaviour and performance is not guaranteed. In the initial state of Mastermind, the empty bottom row contributes to the first guess but has no impact on performance, since all initial guesses yield the same return. This highlights a subtle but important point: a feature can shape the agent’s behaviour without affecting the return that behaviour produces.
The initial state of Mastermind also shows an example of explanations of behaviour and value estimation mirroring each other. In the initial state of Mastermind, the empty bottom row contributes positively to selecting the first guess and negatively to the return estimate. These attributions have the same relative magnitudes but opposite signs, reflecting that the feature increases the likelihood of an action while identifying a low-return state. In contrast, the same features contribute positively to both explanations in the final state: they identify the correct action and a high-return state. This pattern arises because both explanations are defined over the steady-state distribution—they depend on the information a feature provides about the underlying state, which informs both the agent’s behaviour and the return expected from that state.
Both performance and value estimation explain expected return, but differ in how features influence it. Performance explanations attribute responsibility via the agent’s behaviour—what the agent does changes when features are unknown. Value estimation explanations, in contrast, reflect how features inform return estimates without altering behaviour. As shown in the initial state of Mastermind, the bottom row contributes nothing to performance but negatively to value estimation: although it doesn’t change the return received, it signals that the agent is in a low-return state.
8.2 Interpreting SVERL
As with any feature attribution method, interpreting SVERL requires care. While Shapley values reveal how features influence behaviour, performance, or return estimates, they do not necessarily explain why. Human users naturally form beliefs or hypotheses based on these values, but it is easy to overinterpret or misattribute causality. Making these values useful for explanation often depends on insights from psychology and cognitive science, which help bridge the gap between attribution and understanding. SVERL provides a foundation for such interpretation, but further analysis is needed to assess whether the conclusions are well-founded.
Figure 4: Feature contributions to the probability of opening each available square in the two states of Minesweeper shown in Figure 3. The colour of each square shows its Shapley value (scaled between $-1$ and $1$) for contributing to the probability of opening that same square.
Consider the example in Minesweeper. In Figure 3, SVERL assigns negative performance attributions to unopened squares that contain mines in the second state. One might conclude that observing these features lowers performance because it increases the likelihood that the agent opens them. This is a reasonable hypothesis, but it must be validated. Figure 4 shows the contribution of each square to the probability of the agent opening that square. In both states, the unopened squares contribute positively to the probability of being opened. This supports the hypothesis, though it was not a guaranteed conclusion from the original explanation alone.
Misunderstandings can also arise when explaining contributions by focusing too narrowly on individual marginal terms within the Shapley summation. An example of this is the misinterpretation of the $x$ and $y$ contributions in the Taxi domain by Beechey et al. (2023), where these terms were accidentally swapped (corrected in Section B.4). While the qualitative discussion was not entirely incorrect, it misrepresented the full contributions by considering only a subset of feature interactions, overlooking that Shapley values are means over all possible feature coalitions. This highlights the importance of interpreting Shapley-based explanations in light of the full combinatorial structure they represent.
Future work should explore how to effectively support users in interpreting SVERL. This includes developing tools for presenting explanations, educating users on how to interpret feature attributions responsibly, and providing methods for testing and validating the hypotheses they generate. Crucially, these efforts should be grounded in actual use cases and evaluated through user studies to ensure that SVERL contributes meaningfully to understanding and decision-making in practice.
9 Related Work
We now situate the proposed framework within the broader landscape of explainable reinforcement learning. We first review prior work that has applied Shapley values to analyse reinforcement learning agents, before discussing other lines of research in explainability.
9.1 Previous Applications of Shapley Values in Reinforcement Learning
While Shapley values have previously been applied to explain reinforcement learning agents, this work is the first to study them mathematically in this setting, focusing on the validity and meaning of the resulting explanations.
Several prior works analyse the value function using SHAP (Lundberg and Lee, 2017), a Python package that uses marginal characteristic functions to compute Shapley values for models (Zhang et al., 2020; Schreiber et al., 2021; Zhang et al., 2021). Applied to the value function, this setup corresponds to a marginal version of SVERL’s value estimation explanation, which assumes feature independence. However, as shown throughout this paper, value estimation concerns the prediction of return, not how features influence the agent’s behaviour or the return the agent actually receives. Interpreting such explanations as relating to behaviour risks misunderstanding the agent’s interaction with its environment.
Figure 5: Tic-Tac-Toe. Left: feature contributions to the agent’s return estimate (value estimation). Right: contributions to the agent’s performance. The agent (X) plays against a minimax opponent (O) on a $3\times 3$ grid, aiming to place three marks in a row. Players take turns and have an equal probability of starting. Each state is defined by nine features, corresponding to the contents of the grid squares.
This distinction is illustrated in Figure 5, where an optimal agent plays as X against a minimax (Polak, 1989) opponent in Tic-Tac-Toe. In the state shown, the agent must block the opponent’s winning move. The performance explanation correctly attributes responsibility to the opponent’s marked squares. Value estimation, however, assigns zero contribution to all features. This is because the optimal agent consistently forces a draw against the minimax opponent. As a result, the expected return remains fixed at the value corresponding to a draw, independent of the state and, consequently, the values of its features. For an optimal policy, there is no statistical association between feature values and return estimates. While value estimation can identify informative features for prediction, it does not explain behaviour or performance.
Other works use SHAP to explain agent behaviour by analysing the policy (Carbone, 2020; He et al., 2021; Wang et al., 2020; Liessner et al., 2021; Løver et al., 2021; Remman et al., 2022; Theumer et al., 2022). These approaches attribute a deterministic action output by the policy model to individual input features. In continuous domains, this corresponds to SVERL’s continuous behaviour explanation (Section 4.5), where the Gaussian policy’s mean is treated as the selected action and the variance is fixed at zero. In discrete domains, the characteristic function—which averages over actions—can yield non-discrete outputs for unknown features, which do not correspond to any valid action. In contrast, SVERL provides principled behaviour explanations for discrete action spaces by attributing responsibility via action probabilities (Section 4.2).
Another line of work uses SHAP to explain stochastic policies in discrete action spaces by attributing feature importance to action probabilities (Rizzo et al., 2019). This aligns with SVERL’s behaviour explanation using the marginal characteristic function. However, the distribution over which the characteristic function is computed is left unspecified, likely resulting in misinterpretations of the resulting attributions. In contrast, SVERL uses the steady-state distribution to produce principled and interpretable attributions.
Finally, Wäldchen et al. (2022) compute Shapley values by learning characteristic functions through training the policy and the value function with masked features. This approach restricts the number of masked features to stabilise learning, undermining the Shapley principle of evaluating all possible feature subsets. This is an important limitation because a feature’s importance may only emerge when enough others are removed. This limitation prevents the method from satisfying the Shapley axioms and producing principled attributions, unlike SVERL, which is axiomatically sound.
9.2 Other Approaches to Explainable Reinforcement Learning
Many existing approaches to explaining reinforcement learning agents aim to identify important input features or clarify decision behaviour. A widely studied class of methods is based on saliency, where explanations are derived by perturbing inputs and observing changes in outputs (Greydanus et al., 2018; Puri et al., 2020) or by computing input gradients (Wang et al., 2016). While often useful in practice, especially in visual domains, these methods lack theoretical guarantees and can produce inconsistent explanations for different models that yield identical outputs. By contrast, SVERL is grounded in the Shapley framework, ensuring consistency and offering a principled, model-agnostic foundation for attributing importance.
Another prominent direction replaces complex models with simpler, interpretable ones, such as decision trees (Silva et al., 2020; Topin et al., 2021), or trains surrogate models that approximate the behaviour of a trained agent (Bastani et al., 2018; Jhunjhunwala, 2019; Bewley and Lawry, 2021). These methods can improve transparency but often trade off expressivity, with surrogate models failing to represent the original agent faithfully. Other work has sought to better balance interpretability with fidelity by increasing model complexity (Hein et al., 2018; Landajuela et al., 2021), though this inevitably increases cognitive load for users (Dodge et al., 2021). SVERL avoids this trade-off entirely by explaining the behaviour and outcomes of any given agent directly, without retraining or simplification.
Other approaches analyse MDP structure and policies more broadly. Some leverage knowledge of transition dynamics to simulate counterfactuals (Rupprecht et al., 2019), identify causal influences (Madumal et al., 2020), or estimate task completion probabilities (Cruz et al., 2019, 2023). Others rely on assumptions about reward decomposability to trace how different objectives contribute to value estimates (Juozapaitis et al., 2019) or identify critical states where incorrect decisions would lead to poor performance (Amir et al., 2018; Huang et al., 2018). These methods can offer useful insights but typically depend on explicit assumptions about environment models or reward structure, or are limited in scope, highlighting states without explaining why actions are chosen. By contrast, SVERL provides general-purpose, instance-level behaviour explanations without making such assumptions.
10 Limitations and Future Work
This paper has focused on establishing the theoretical foundations of SVERL as a unified framework for explaining various elements of reinforcement learning agents. There is significant scope for extension. In this section, we highlight several promising directions that build on the foundations we have presented here, from practical challenges to broader theoretical developments.
10.1 Explaining Agent-Environment Interactions More Broadly
We have focused on explaining agent behaviour at specific states or for specific actions. Understanding behaviour broadly often involves reasoning about trends across many interactions. While one could present the Shapley values for many individual states and actions, doing so imposes a high cognitive load and hinders comprehension (Covert et al., 2021). Here, we explore how SVERL could be extended to summarise explanations across states or actions to provide broader insights.
A natural idea is to aggregate SVERL’s explanations for individual states or actions by taking their expectation under some distribution $\mathcal{D}$. For example, given a behaviour explanation $\phi_{i}(\tilde{\pi}_{s}^{a})$ for action $a$ at state $s$, we compute:
$\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}[\phi_{i}(\tilde{\pi}_{s}^{a})],$ (12)
as a general framework for summarising feature contributions. While we illustrate this approach for explanations of behaviour, the same approach applies to performance and value estimation. The choice of distribution $\mathcal{D}$ determines the kind of insight obtained. It could reflect, for example, the full range of the agent’s experience, focus on particularly salient states, or highlight trends conditional on specific feature values.
Under this framework, one might explore many directions. We now consider two natural first ideas: aggregating explanations across states using the steady-state distribution and aggregating across actions using the policy. Each may offer different perspectives on how features influence behaviour or outcomes broadly.
Averaging over the steady-state distribution.
Let $\mathcal{D}=p^{\pi}(s)$, the steady-state distribution under policy $\pi$. Then, for behaviour explanations, we compute:
$\underset{s\sim p^{\pi}(s)}{\mathbb{E}}[\phi_{i}(\tilde{\pi}_{s}^{a})],$
which represents the average contribution of feature $i$ to the probability of action $a$ across the agent’s experience, weighting each feature contribution by the likelihood it’s observed. However, perhaps counterintuitively, this quantity is always zero.
Proposition 21
For any policy $\pi$, the expected contribution of any feature to the probability of action $a$ under the steady-state distribution is zero:
$\underset{s\sim p^{\pi}(s)}{\mathbb{E}}[\phi_{i}(\tilde{\pi}_{s}^{a})]=0.$
This result follows from the efficiency axiom: Shapley values explain deviations from a baseline, which is defined as the expected action probability over $p^{\pi}(s)$. On average, these deviations must sum to zero over $p^{\pi}(s)$. A similar argument applies to value estimation, as stated below.
Corollary 22
For any policy $\pi$, the expected contribution of any feature to the return estimate under the steady-state distribution is zero:
$\underset{s\sim p^{\pi}(s)}{\mathbb{E}}[\phi_{i}(\hat{v}^{\pi}_{s})]=0.$
In both cases, while local explanations offer insight into individual decisions or predictions, their steady-state averages provide no signal. This does not rule out the use of other state distributions. Alternative choices for $\mathcal{D}$ may yield meaningful summaries, making this a promising direction for future work.
In contrast, performance explanations can yield non-zero expectations:
$\underset{s\sim p^{\pi}(s)}{\mathbb{E}}[\phi_{i}(\tilde{v}^{\pi}_{s})]\neq 0.$
Here, the baseline is not defined as an expectation over $p^{\pi}(s)$, so feature attributions retain signal when averaged. However, whether this expectation provides useful insight remains unclear. We leave a full analysis of its interpretability to future work.
Averaging over the policy distribution.
As another example, let $\mathcal{D}=\pi(s,a)$, the policy’s action distribution at a given state. Then, for behaviour explanations, we compute:
$\underset{a\sim\pi(s,a)}{\mathbb{E}}[\phi_{i}(\tilde{\pi}_{s}^{a})],$
which represents the average contribution of feature $i$ to the probabilities of actions in state $s$, weighting each feature contribution by the likelihood that the action is chosen. One could interpret this concept as capturing the general influence of a feature value on behaviour across actions at a given state. By the linearity of Shapley values, this quantity is itself a Shapley value for a new coalitional game with characteristic function defined as the expected action probability:
$\underset{a\sim\pi(s,a)}{\mathbb{E}}[\phi_{i}(\tilde{\pi}_{s}^{a})]=\phi_{i}\left(\underset{a\sim\pi(s,a)}{\mathbb{E}}[\tilde{\pi}_{s}^{a}]\right).$
Its interpretation remains unclear, but this connection to cooperative game theory offers a promising direction for investigating what insight such a summary explanation provides, and whether it meaningfully captures a feature’s importance across the agent’s behaviour.
10.2 Approximating SVERL
Computing SVERL’s explanations is computationally demanding. Each characteristic function involves an expectation over the entire state space, and each Shapley value is defined as a sum over all $2^{|{\cal F}|}$ subsets of features. These requirements quickly become intractable as the number of features or states grows. As a result, practical applications of SVERL in large domains require approximation techniques.
Approximation techniques have been widely studied in the supervised learning setting, and it is only natural to expect that many of these methods could be extended to SVERL. Here, we adapt one such method, the Monte Carlo sampling techniques in Section 2.5, to approximate SVERL’s explanations of behaviour, performance, and value estimation. We present this theoretical outline to demonstrate that SVERL can be approximated in practice, leaving its full treatment and discussion of alternative methods to future work.
We begin with explanations of behaviour and value estimation. We propose approximating their characteristic functions, defined in Equations 8 and 11, using Monte Carlo sampling:
$\tilde{\pi}_{s}^{a}(\mathcal{C})$ $=\mathbb{E}[\pi(S,a)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\lim_{n\to\infty}\frac{1}{n}\sum_{s^{\prime}\sim p^{\pi}(\cdot\mid S_{\mathcal{C}}=s_{\mathcal{C}})}\pi(s^{\prime},a),$ (13)
$\hat{v}^{\pi}_{s}(\mathcal{C})$ $=\mathbb{E}[\hat{v}^{\pi}(S)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\lim_{n\to\infty}\frac{1}{n}\sum_{s^{\prime}\sim p^{\pi}(\cdot\mid S_{\mathcal{C}}=s_{\mathcal{C}})}\hat{v}^{\pi}(s^{\prime}).$ (14)
Paralleling the approach in supervised learning, we combine these approximations with the permutation-based Shapley value definition in Equation 2, yielding a single Monte Carlo sampling method:
$\phi_{i}(\tilde{\pi}_{s}^{a})$ $=\lim_{n\to\infty}\frac{1}{n}\sum_{\begin{subarray}{c}s^{\prime}\sim p^{\pi}(\cdot\mid S_{\mathcal{C}\cup\{i\}}=s_{\mathcal{C}\cup\{i\}})\\ s^{\prime\prime}\sim p^{\pi}(\cdot\mid S_{\mathcal{C}}=s_{\mathcal{C}})\end{subarray}}\left[\pi(s^{\prime},a)-\pi(s^{\prime\prime},a)\right],$
$\phi_{i}(\hat{v}^{\pi}_{s})$ $=\lim_{n\to\infty}\frac{1}{n}\sum_{\begin{subarray}{c}s^{\prime}\sim p^{\pi}(\cdot\mid S_{\mathcal{C}\cup\{i\}}=s_{\mathcal{C}\cup\{i\}})\\ s^{\prime\prime}\sim p^{\pi}(\cdot\mid S_{\mathcal{C}}=s_{\mathcal{C}})\end{subarray}}\left[\hat{v}^{\pi}(s^{\prime})-\hat{v}^{\pi}(s^{\prime\prime})\right],$
where $\mathcal{C}$ is the set of features that precede $i$ in ordering $O$ and $O$ is sampled uniformly from $\pi({\cal F})$, the set of all permutations of ${\cal F}$.
Performance explanations require a different treatment. The characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$, defined in Equation 10, is the expected return from state $s$ under a policy that follows the partially observed policy $\tilde{\pi}_{s}^{a}(\mathcal{C})$ in $s$, and the fully observed policy $\pi$ elsewhere.
Because this characteristic is an expected return, it can be estimated using any standard reinforcement learning method. However, it depends on the policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$, which must also be approximated. We propose incorporating the sampling-based approximation of $\tilde{\pi}_{s}^{a}(\mathcal{C})$ from Equation 13 directly into the policy definition used within the performance characteristic’s expectation:
$\tilde{v}^{\pi}_{s}(\mathcal{C})$ $=\mathbb{E}_{\hat{\pi}_{\mathcal{C}}}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\mid S_{t}=s\right],$
$\text{where}\quad\hat{\pi}_{\mathcal{C}}(a_{t}\mid s_{t})$ =$=\begin{cases}\pi(a_{t}\mid s^{\prime}),&\text{with }s^{\prime}\sim p^{\pi}(\cdot\mid s_{\mathcal{C}})\quad\text{if }s_{t}=s,\\ \pi(a_{t}\mid s_{t}),&\text{otherwise.}\end{cases}$
This defines a new policy $\hat{\pi}_{\mathcal{C}}(a_{t}\mid s_{t})$ that, when visiting $s$, selects actions based on a sampled state $s^{\prime}$ consistent with the observed features: $s_{\mathcal{C}}=s^{\prime}_{\mathcal{C}}$, and follows the fully observed policy $\pi$ elsewhere. By rewriting the performance characteristic in this equivalent form, we embed the approximation of the policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$ into the return estimation itself, allowing $\tilde{v}^{\pi}_{s}(\mathcal{C})$ to be computed directly using standard reinforcement learning methods.
Finally, with $\tilde{v}^{\pi}_{s}(\mathcal{C})$ approximated in this way, the Shapley values for performance can be computed using the usual permutation-based sum:
$\phi_{i}(\tilde{v}^{\pi}_{s})=\lim_{n\to\infty}\frac{1}{n}\sum_{O\sim\pi({\cal F})}\left[\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})-\tilde{v}^{\pi}_{s}(\mathcal{C})\right].$
All three explanations—behaviour, performance, and value estimation—require access to the conditional steady-state distribution $p^{\pi}(S\mid S_{\mathcal{C}}=s_{\mathcal{C}})$. This distribution must be learned in large domains, presenting an important direction for future work. One promising starting point is the parametric approach introduced by Frye et al. (2020a), which uses variational inference to learn conditional data distributions from a dataset. A natural adaptation to the reinforcement learning setting would be to use this method to learn the conditional steady-state distribution from a replay buffer of agent experience.
Here we have presented an outline for approximating SVERL by adapting Monte Carlo techniques from supervised learning to the reinforcement learning setting, without investigating empirical performance or efficiency. We encourage a full treatment of these questions in subsequent research. Exploring other approximation methods developed in supervised learning may also prove fruitful. For example, Frye et al. (2020a) propose learning parametric models of the characteristic function directly, and Jethani et al. (2021) amortises Shapley value approximations across inputs by training parametric models to predict them. These approaches are promising future avenues to enable more scalable or efficient approximations of SVERL.
10.3 Using Value Estimation to Inform Behaviour
SVERL’s value estimation explanations reveal how feature values influence the agent’s return estimate. As shown in the examples in Section 7, feature values in a given state that also appear in high-return states tend to receive positive attributions. In contrast, feature values associated with lower-return states tend to receive negative attributions. This suggests that these explanations are useful for highlighting which aspects of the environment correlate with successful estimated outcomes. One intriguing possibility is to use this information to inform behaviour—for example, encouraging actions that lead to positively attributed features, or identifying negatively attributed features that should be changed to improve expected return. The additional example in Section B.3 illustrates this idea concretely: in a dice-rolling domain, each die is treated as a feature and receives a negative attribution if and only if the optimal policy re-rolls the corresponding die to change its value. How to effectively incorporate these signals into decision-making or policy optimisation is a compelling question, and one we believe represents an important direction for future work.
10.4 Connecting SVERL to Broader Work on Shapley Values
SVERL provides a principled framework for explaining reinforcement learning agents using Shapley values. It introduces new cooperative games tailored to reinforcement learning by defining policy, performance, and value estimation characteristic functions. Because the framework inherits the structure and axiomatic foundations of classic Shapley value theory, it can naturally accommodate alternative definitions, theoretical developments, and methodological advances from the broader literature.
To illustrate the kind of opportunities this generality creates, we highlight two possible extensions. First, Frye et al. (2020b) propose computing Shapley values under a causal model, modifying the Shapley value sum to account for known dependencies between features. Extending SVERL in this way could yield explanations that more accurately reflect the causal structure of the environment or agent behaviour. Another possibility is to explore alternative solution concepts from cooperative game theory. For instance, Banzhaf values (Banzhaf III, 1964) define a different notion of attribution based on a distinct set of axioms and may offer complementary insights when applied within the SVERL framework. These examples demonstrate only two possible ways broader theoretical developments in Shapley value research can be incorporated into reinforcement learning using SVERL. Exploring these connections provides a promising avenue for future work.
11 Conclusion
Reinforcement learning has achieved impressive results across many domains, often surpassing human ability. Yet the how and why of agents interacting with their environments remains difficult to interpret, which hinders trust and limits adoption. In this work, we addressed this challenge by introducing Shapley Values for Explaining Reinforcement Learning (SVERL), a theoretical framework for explaining agent-environment interactions.
Motivated by an illustrative example, we identified three core elements of reinforcement learning to explain: behaviour, performance, and value estimation. Together, these offer a general understanding of reinforcement learning systems. We proposed a unified framework for explaining these elements through the influence of features at specific states. By treating features as players that cooperate to produce each element, we defined a family of theoretically principled explanations grounded in the well-established foundations of Shapley values in cooperative game theory.
We examined how these explanations relate to one another and how the framework as a whole can provide a more complete view of agent–environment interactions. We also discussed important considerations for interpreting SVERL in practice, including potential challenges and limitations. Finally, we considered possible extensions of the framework, such as explaining behaviour more broadly and approximating SVERL in large-scale settings.
By providing a theoretical foundation for explaining reinforcement learning, we aim to support the development of principled and interpretable reinforcement learning methods. We hope that this work lays the groundwork for future research towards more transparent and trustworthy systems.
Acknowledgments and Disclosure of Funding
This work was supported by the UKRI Centre for Doctoral Training in Accountable, Responsible and Transparent AI (ART-AI) [EP/S023437/1], the EPSRC Centre for Doctoral Training in Digital Entertainment (CDE) [EP/L016540/1] and the University of Bath. This research made use of Hex, the GPU Cloud in the Department of Computer Science at the University of Bath.
Appendix A Proofs of Propositions and Results
We present each result and its associated proof in full detail.
Proposition 2
The Shapley values $\phi_{1}(\tilde{\pi}_{s}^{a}),\dots,\phi_{n}(\tilde{\pi}_{s}^{a})$ for the discrete policy game $({\cal F},\tilde{\pi}_{s}^{a})$ are the unique allocation of the action probability $\pi(s,a)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\tilde{\pi}_{s}^{a}:2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the policy game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\tilde{\pi}_{s}^{a}),\dots,\phi_{n}(\tilde{\pi}_{s}^{a})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
Proposition 4
When comparing policies using mean squared error between the probabilities of selecting action $a$, the policy $\mu$ is the policy that deviates the least in expectation from the policy $\pi$:
$\mu=\operatorname*{arg\,min}_{g}\mathbb{E}\left[\lVert\pi(S,a)-g(S_{\mathcal{C}},a)\rVert^{2}\right],$ (15)
where $g:{\cal S}_{\mathcal{C}}\rightarrow\Delta({\cal A})$ is also defined over the subset of features $\mathcal{C}$.
Proof  We rearrange the expected squared error:
$\mathbb{E}\left[\lVert\pi(S,a)-g(S_{\mathcal{C}},a)\rVert^{2}\right]$ $=\mathbb{E}_{S}\left[\lVert\pi(S,a)-\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$\quad+\mathbb{E}_{S_{\mathcal{C}}}\left[\lVert\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]-g(S_{\mathcal{C}},a)\rVert^{2}\right].$
The first term is independent of $g$, and the second is minimised when $g(S_{\mathcal{C}},a)=\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]=\mu(S_{\mathcal{C}},a)$. Substituting $g=\mu$ gives:
$\mathbb{E}\left[\lVert\pi(S,a)-\mu(S_{\mathcal{C}},a)\rVert^{2}\right]$ $=\mathbb{E}_{S}\left[\lVert\pi(S,a)-\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$\quad+\mathbb{E}_{S_{\mathcal{C}}}\left[\lVert\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]-\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$=\mathbb{E}_{S}\left[\lVert\pi(S,a)-\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}]\rVert^{2}\right],$
confirming that $\mu$ achieves the minimum.
Proposition 8
The Shapley values $\phi_{1}(\tilde{\mu}_{s}),\dots,\phi_{n}(\tilde{\mu}_{s})$ for the continuous policy game $({\cal F},\tilde{\mu}_{s})$ are the unique allocation of the expected action $\mu(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\tilde{\mu}_{s}:2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the continuous policy game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\tilde{\mu}_{s}),\dots,\phi_{n}(\tilde{\mu}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
Proposition 10
When comparing mean functions using mean squared error, the function $\omega$ deviates the least in expectation from the mean $\mu$ of the full policy $\pi$:
$\omega=\operatorname*{arg\,min}_{g:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}}\mathbb{E}\left[\lVert\mu(S)-g(S_{\mathcal{C}})\rVert^{2}\right].$ (16)
Proof  We rearrange the expected squared error:
$\mathbb{E}\left[\lVert\mu(S)-g(S_{\mathcal{C}})\rVert^{2}\right]$ $=\mathbb{E}_{S}\left[\lVert\mu(S)-\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$\quad+\mathbb{E}_{S_{\mathcal{C}}}\left[\lVert\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]-g(S_{\mathcal{C}})\rVert^{2}\right].$
The first term is independent of $g$, and the second is minimised when $g(S_{\mathcal{C}})=\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]=\omega(S_{\mathcal{C}})$. Substituting $g=\omega$ gives:
$\mathbb{E}\left[\lVert\mu(S)-\omega(S_{\mathcal{C}})\rVert^{2}\right]$ $=\mathbb{E}_{S}\left[\lVert\mu(S)-\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$\quad+\mathbb{E}_{S_{\mathcal{C}}}\left[\lVert\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]-\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$=\mathbb{E}_{S}\left[\lVert\mu(S)-\mathbb{E}[\mu(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right],$
confirming that $\omega$ achieves the minimum.
Proposition 13
The Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ for the performance game $({\cal F},\tilde{v}^{\pi}_{s})$ are the unique allocation of the expected return $v^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C}):2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the performance game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
Proposition 17
The Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ for the value estimation game $({\cal F},\hat{v}^{\pi}_{s})$ are the unique allocation of the estimated return $\hat{v}^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\hat{v}^{\pi}_{s}:2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the value estimation game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
Proposition 19
When comparing return estimates using mean squared error, the function $\hat{u}^{\pi}$ is the closest in expectation to $\hat{v}^{\pi}$ among all functions defined over the subset of features $\mathcal{C}$:
$\hat{u}^{\pi}=\operatorname*{arg\,min}_{g:{\cal S}_{\mathcal{C}}\rightarrow\mathbb{R}}\mathbb{E}\left[\lVert\hat{v}^{\pi}(S)-g(S_{\mathcal{C}})\rVert^{2}\right].$
Proof  We rearrange the expected squared error:
$\mathbb{E}\left[\lVert\hat{v}^{\pi}(S)-g(S_{\mathcal{C}})\rVert^{2}\right]$ $=\mathbb{E}_{S}\left[\lVert\hat{v}^{\pi}(S)-\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$\quad+\mathbb{E}_{S_{\mathcal{C}}}\left[\lVert\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]-g(S_{\mathcal{C}})\rVert^{2}\right].$
The first term is independent of $g$, and the second is minimised when $g(S_{\mathcal{C}})=\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]=\hat{u}^{\pi}(S_{\mathcal{C}})$. Substituting $g=\hat{u}^{\pi}$ gives:
$\mathbb{E}\left[\lVert\hat{v}^{\pi}(S)-\hat{u}^{\pi}(S_{\mathcal{C}})\rVert^{2}\right]$ $=\mathbb{E}_{S}\left[\lVert\hat{v}^{\pi}(S)-\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$\quad+\mathbb{E}_{S_{\mathcal{C}}}\left[\lVert\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]-\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right]$
$=\mathbb{E}_{S}\left[\lVert\hat{v}^{\pi}(S)-\mathbb{E}[\hat{v}^{\pi}(S)\,|\,S_{\mathcal{C}}]\rVert^{2}\right],$
confirming that $\hat{u}^{\pi}$ achieves the minimum.
Proposition 21
For any policy $\pi$, the expected contribution of any feature to the probability of action $a$ under the steady-state distribution is zero:
$\underset{s\sim p^{\pi}(s)}{\mathbb{E}}[\phi_{i}(\tilde{\pi}_{s}^{a})]=0.$
Proof  Begin by expanding the definition of the Shapley value and applying the linearity of expectation:
$\mathbb{E}_{p^{\pi}(s)}[\phi_{i}(\tilde{\pi}_{s}^{a})]$ $=\mathbb{E}_{p^{\pi}(s)}\left[\sum_{\mathcal{C}\subseteq{\cal F}\setminus\{i\}}\frac{|\mathcal{C}|!(|{\cal F}|-|\mathcal{C}|-1)!}{|{\cal F}|!}\left[\tilde{\pi}_{s}^{a}(\mathcal{C}\cup\{i\})-\tilde{\pi}_{s}^{a}(\mathcal{C})\right]\right]$
$=\sum_{\mathcal{C}\subseteq{\cal F}\setminus\{i\}}\frac{|\mathcal{C}|!(|{\cal F}|-|\mathcal{C}|-1)!}{|{\cal F}|!}\left[\mathbb{E}_{p^{\pi}(s)}[\tilde{\pi}_{s}^{a}(\mathcal{C}\cup\{i\})]-\mathbb{E}_{p^{\pi}(s)}[\tilde{\pi}_{s}^{a}(\mathcal{C})]\right].$
To simplify each expectation term, recall that $\tilde{\pi}_{s}^{a}(\mathcal{C})$ is defined as the conditional expectation of $\pi(s,a)$ given $S_{\mathcal{C}}=s_{\mathcal{C}}$, followed by marginalisation over $s_{\mathcal{C}}$:
$\mathbb{E}_{p^{\pi}(s)}[\tilde{\pi}_{s}^{a}(\mathcal{C})]$ $=\mathbb{E}_{s_{\mathcal{C}}\sim p^{\pi}(s_{\mathcal{C}})}\left[\mathbb{E}[\pi(S,a)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]\right]=\mathbb{E}_{p^{\pi}(s)}[\pi(S,a)].$
That is, marginalising the conditional expectation recovers the unconditional expectation. The same holds for $\mathcal{C}\cup\{i\}$, so:
$\mathbb{E}_{p^{\pi}(s)}[\tilde{\pi}_{s}^{a}(\mathcal{C}\cup\{i\})]=\mathbb{E}_{p^{\pi}(s)}[\pi(S,a)].$
Substituting back, the two terms in the Shapley difference cancel:
$\mathbb{E}_{p^{\pi}(s)}[\phi_{i}(\tilde{\pi}_{s}^{a})]$ $=\sum_{\mathcal{C}\subseteq{\cal F}\setminus\{i\}}\frac{|\mathcal{C}|!(|{\cal F}|-|\mathcal{C}|-1)!}{|{\cal F}|!}\left[\mathbb{E}[\pi(S,a)]-\mathbb{E}[\pi(S,a)]\right]$
$=0.$
This concludes the proof.
Appendix B Additional Examples
Additional examples are presented. We use three intuitive examples to illustrate how our framework attributes behaviour, performance and value estimates to features. We then conclude with an additional example explaining behaviour, performance, and value estimation.
B.1 Additional Example on Explaining Behaviour
State ($s$) Action ($a$) $p^{\pi}(s)$ $\tilde{\pi}_{s}^{a}(\text{Both})$ $\tilde{\pi}_{s}^{a}(\text{I})$ $\tilde{\pi}_{s}^{a}(\text{C})$ $\tilde{\pi}_{s}^{a}(\emptyset)$ $\phi_{\text{I}}$ $\phi_{\text{C}}$ $(1,\text{Red})$ N $0.25$ 00 00 00 $0.25$ $-0.125$ $-0.125$ E $1$ $1$ $1$ $0.25$ $0.375$ $0.375$ S 00 00 00 $0.25$ $-0.125$ $-0.125$ W 00 00 00 $0.25$ $-0.125$ $-0.125$ $(3,\text{Green})$ N $0.25$ $1$ $1$ $0.5$ $0.25$ $0.625$ $0.125$ E 00 00 00 $0.25$ $-0.125$ $-0.125$ S 00 00 00 $0.25$ $-0.125$ $-0.125$ W 00 00 $0.5$ $0.25$ $-0.375$ $0.125$
[…]
We first consider explaining agent behaviour. Figure 6 shows a $2\times 2$ gridworld in which each state is defined by two features: its index and its colour. For example, the top-left state is indexed as 1, coloured red, and denoted $(1,\text{Red})$. The figure also shows a policy $\pi$ that moves clockwise through the grid: East in state $(1,\text{Red})$, South in state $(2,\text{Blue})$, West in state $(4,\text{Green})$, and North in state $(3,\text{Green})$. We analyse how the agent’s action probabilities in states $(1,\text{Red})$ and $(3,\text{Green})$ change as different features become known. These probabilities define SVERL’s characteristic values for explaining behaviour, from which the Shapley values are computed. Both are summarised in the table in Figure 6.
Consider first state $(1,\text{Red})$. When neither feature is known, the agent is equally likely to be in any state (by the steady-state distribution), and thus acts uniformly over the four actions. If the index becomes known to be $1$, or the colour becomes known to be Red, the agent must be in state $(1,\text{Red})$ and selects East with probability 1. Thus, both features fully resolve the agent’s uncertainty and have the same effect on the action probabilities. This is reflected in the Shapley values in Figure 6, which assign equal contributions to the index and colour for increasing the probability of the correct action and decreasing the probabilities of the incorrect ones.
Now consider state $(3,\text{Green})$. If the index becomes known to be $3$, the agent must be in state $(3,\text{Green})$ and selects North with probability 1. If instead the colour becomes known to be Green, the agent could be in either state $(3,\text{Green})$ or $(4,\text{Green})$, and hence selects North and West with equal probability. Thus, both features increase the probability of the correct action (North), the index increases this probability more, and the colour increases the probability of an incorrect action (West). This distinction is reflected in the Shapley values in Figure 6.
B.2 Additional Example on Explaining Performance
1
2
3
5
4
G
G
$x$
$y$
State $(s)$ Action ($a$) $\tilde{\pi}_{s}^{a}(\text{Both})$ $\tilde{\pi}_{s}^{a}(x)$ $\tilde{\pi}_{s}^{a}(y)$ $\tilde{\pi}_{s}^{a}(\emptyset)$
1 N 0 0 $2/3$ $6/7$
E 1 1 $1/3$ $1/7$
S 0 0 0 0
W 0 0 0 0
2 N 1 1 $2/3$ $6/7$
E 0 0 $1/3$ $1/7$
S 0 0 0 0
W 0 0 0 0
State $(s)$ $p^{\pi}(s)$ $\tilde{v}^{\pi}_{s}(\text{Both})$ $\tilde{v}^{\pi}_{s}(x)$ $\tilde{v}^{\pi}_{s}(y)$ $\tilde{v}^{\pi}_{s}(\emptyset)$ $\phi_{x}$ $\phi_{y}$
1 $1/7$ $6.00$ $6.00$ $4.00$ $0.00$ $4.00$ $2.00$
2 $2/7$ $7.00$ $7.00$ $6.50$ $6.83$ $0.33$ $-0.17$
Figure 8: A five-state gridworld with two goal states (G). The agent starts in state 1 or 2 with equal probability and follows the optimal deterministic policy (shown with arrows). Each state is defined by two features: its $x$- and $y$-coordinates. The top-right table shows the action probabilities under different feature subsets for states 1 and 2. The bottom table shows the characteristic values and Shapley attributions used to explain performance.
We now consider explaining performance. Figure 8 shows a five-state gridworld where the agent begins uniformly at random in state 1 or 2, receiving a reward of $-1$ per action and an additional $+10$ upon reaching a goal state. The figure also shows the optimal deterministic policy $\pi$: East from state 1 and North from all others; invalid actions incur reward but leave the agent’s position unchanged. Each state is described by two features: its $x$- and $y$-coordinates. We analyse how the expected return from each state changes as these coordinates become known, owing to the agent’s behaviour in that state shifting accordingly. These returns define SVERL’s characteristic values for explaining performance, while the resulting Shapley values quantify the influence of each feature. Both are shown in the bottom table of Figure 8; the top-right table shows the corresponding action probabilities under different subsets of known features.
Consider state 1. When neither feature is known, the agent has probability $1/7$ of being in state 1 (with optimal action East) and probability $2/7$ of being in states 2, 3, and 4 (all with optimal action North). Therefore, the agent selects North with probability $6/7$ and East with $1/7$, yielding an expected return of 00 from state 1. If the $x$-coordinate becomes known, the agent must be in state 1 and selects East, yielding a return of $6$. If instead the $y$-coordinate becomes known, the agent could be in state 1 or 2, with state 2 twice as likely, and so selects North with probability $2/3$ and East with $1/3$, giving an expected return of $4$. This difference is reflected in the Shapley values in Figure 8, which assign twice as much influence on expected return to the $x$-coordinate as the $y$-coordinate.
Now consider state 2. As in state 1, when no features are known, the agent follows the steady-state distribution, selecting North with probability $6/7$ and East with probability $1/7$, yielding an expected return of $6.83$ from state 2. If the $x$-coordinate becomes known, the agent must be in states 2, 3, or 4 (all with optimal action North). Therefore, the agent deterministically selects North, increasing the return from state 2 to $7$. If instead the $y$-coordinate becomes known, the agent could be in state 1 or 2, with state 2 twice as likely, and so selects North with probability $2/3$ and East with $1/3$. This gives a lower expected return of $6.5$ because the likelihood of the optimal action (North) has decreased from $6/7$ to $2/3$. The Shapley values in Figure 8 reflect this: the $x$-coordinate contributes positively to performance, while the $y$-coordinate has a negative influence.
B.3 Additional Example on Explaining Value Estimation
Table 7: Characteristic values and Shapley values for value estimates in two states of a dice-rolling game. Each state is defined by two features: the values of two six-sided dice, denoted as $(d_{1},d_{2})$. The agent may re-roll either die, both dice, or neither. After each action, the episode ends with probability $0.5$, yielding a reward of $1$ if the sum of the dice exceeds $9$, and 00 otherwise. For the two states shown, the optimal policy re-rolls the $3$ in $(3,6)$ and re-rolls both dice in $(1,1)$.
State $(s)$ $p^{\pi}(s)$ $\hat{v}^{\pi}_{s}(\text{Both})$ $\hat{v}^{\pi}_{s}(d_{1})$ $\hat{v}^{\pi}_{s}(d_{2})$ $\hat{v}^{\pi}_{s}(\emptyset)$ $\phi_{d_{1}}$ $\phi_{d_{2}}$
$(3,6)$ $0.024$ $0.67$ $0.45$ $0.90$ $0.66$ $-0.22$ $0.23$
$(1,1)$ $0.018$ $0.36$ $0.45$ $0.45$ $0.66$ $-0.15$ $-0.15$
Next, we consider explaining value estimation in a dice-rolling game. Each episode begins with a random roll of two six-sided dice. The resulting state is denoted $(d_{1},d_{2})$, where $d_{1}$ and $d_{2}$ are the first and second die values, respectively. The agent may re-roll either die, both dice, or neither. After each action, the episode ends with probability $0.5$, yielding a reward of $1$ if the sum of the dice is $10$ or more, and 00 otherwise. The optimal policy keeps $5$s and $6$s, and re-rolls $1$s, $2$s, $3$s, and $4$s. We analyse how the agent’s estimated return in each state changes as one or both dice become known. These returns define the characteristic values in SVERL’s value estimation game. The resulting Shapley values quantify each die’s contribution, as shown in Table 7.
Consider state $(3,6)$. When neither die is known, the value estimate is the average return over the steady-state distribution: $0.66$. If $d_{2}$ becomes known to be $6$, the estimate increases to $0.90$ because states with a $6$ have higher average returns—the agent keeps the $6$ and only possibly re-rolls the other die. If instead $d_{1}$ becomes known to be $3$, the estimate drops to $0.45$ because states with a $3$, which are re-rolled by the optimal policy, have lower average returns. The Shapley values reflect this difference: the $6$ contributes positively to the value estimate, while the $3$ contributes negatively.
Now consider state $(1,1)$. As above, when neither die is known, the estimate is $0.66$. If either die becomes known to be $1$, the estimate decreases to $0.45$ because states with a $1$ have a lower average return—the $1$ must be re-rolled. When both dice are known, the agent must be in the lowest-return state, and the estimate falls further to $0.36$. Thus, each die lowers the value estimate by the same amount, as reflected in their equal negative Shapley values in Table 7.
Figure 9: Shapley values for the value estimates of every state in the dice-rolling game. Each point corresponds to a state, with the horizontal and vertical axes showing the Shapley value for the first and second die, respectively. Points are coloured according to which dice the optimal policy re-rolls. The quadrant structure perfectly aligns with negative contributions and the need to re-roll, suggesting that value-based attributions can signal which features should be altered to improve return.
Interestingly, this pattern holds across the entire state space. Figure 9 shows the Shapley values for each die in every state, with each point coloured according to which dice are re-rolled under the optimal policy. The quadrant structure reveals a perfect correspondence: a die receives a negative attribution if and only if it is re-rolled. This illustrates how value-based explanations could inform control, highlighting features that should be changed to improve return—an idea we visited in our future work discussion (Section 10.3).
B.4 Additional Example on Explaining Behaviour, Performance and Value Estimation
Refer to caption Refer to caption
Figure 10: Shapley value attributions for two states in Taxi (top and bottom). First column: feature contributions to the probability of selecting the optimal action taken (behaviour). Second column: contributions to the agent’s performance. Third column: contributions to the agent’s return estimate (value estimation). Fourth column: visualisation of the corresponding state. The taxi’s location is marked with a rectangle, the passenger location with a p, and the destination is circled.
Finally, we consider explaining policy, performance and value estimation in the classic Taxi domain (Dietterich, 2000). In this domain, an agent must navigate a gridworld to pick up a passenger and deliver them to their destination. Each state is described by four features: the taxi’s $(x,y)$ position, the passenger location (R, G, B, Y, or in-taxi), and the destination (R, G, B, or Y). Each step incurs a reward of $-1$, with a bonus of $+20$ for a successful drop-off and a penalty of $-10$ for inappropriate pick-up and drop-off actions. We use the OpenAI Gym Taxi-v3 implementation (Brockman et al., 2016); see there for further details. Using our proposed framework, we analyse two representative states from an optimal policy (Figure 10).
In the top state, the taxi moves south towards the passenger at location B, and the destination is location G. The behaviour explanation highlights the passenger location as the dominant contributor, with the $x$ and $y$ coordinates contributing similarly. Surprisingly, the destination receives a small negative attribution, likely because knowing it without the passenger’s location can increase the chance of taking suboptimal actions. For performance, the $x$ and $y$ coordinates contribute positively, though $x$ more than $y$, illustrating that while these features are equally important for selecting actions, their consequences for return may differ. For value estimation, the $y$ coordinate and passenger location lower the return estimate, reflecting that the agent is not in the same square as the passenger and has not picked them up.
In the bottom state, the taxi moves south to deliver the passenger, now onboard, to the destination at location B. The behaviour explanation now highlights the passenger and destination locations as similar key contributors. This illustrates the agent’s competing goals: keep the passenger in the taxi and drop them off at the destination. The performance explanation mirrors this pattern. In contrast to the top state, the passenger’s presence in the taxi—indicating a higher return—now increases the return estimate for value estimation.
This example shows how the same features influence each explanation differently depending on their role in the task and the agent’s progress toward its objective.